### v1

# Meta Pruning via Graph Metanetworks : A Meta Learning Framework for Network Pruning

Abstract
Network pruning, aimed at reducing network size while preserving accuracy, has attracted significant research interest. Numerous pruning techniques have been proposed over time. They are becoming increasingly effective, but more complex and harder to interpret as well. Given the inherent complexity of neural networks, we argue that manually designing pruning criteria has reached a bottleneck. To address this, we propose a novel approach in which we use a neural network to prune neural networks. More specifically, we introduce the newly developed idea of metanetwork from meta-learning into pruning. A metanetwork is a network that takes another network as input and produces a modified network as output. In this paper, we first establish a bijective mapping between neural networks and graphs, and then employ a graph neural network as our metanetwork. We train a metanetwork that learns the pruning strategy automatically which can transform a network that is hard to prune into another network that is much easier to prune. Once the metanetwork is trained, our pruning needs nothing more than a feedforward through the metanetwork and the standard finetuning to prune at state-of-the-art. Our method achieved outstanding results on many popular and representative pruning tasks (including ResNet56 on CIFAR10, VGG19 on CIFAR100, ResNet50 on ImageNet). Our code is available at https://github.com/Yewei-Liu/MetaPruning.
1 Introduction
Correspondence to Muhan Zhang
With the rapid advancement of deep learning [33, 53], neural networks have become increasingly powerful. However, this improved performance often comes with a significant increase in the number of parameters and computational cost (FLOPs). As a result, there is growing interest in methods to simplify these networks while preserving their performance. Pruning, which involves selectively removing certain parts of a neural network, has proven to be an effective approach [2, 17, 51].
A key idea of a large number of previous pruning works is to remove the unimportant components of a neural network. To achieve this goal, various criteria that measure the importance of components of a neural network have been invented. Some criteria are based on norm or magnitude scores [15, 36, 55, 18]. Some are based on how much the performance changes when certain parts of the network are removed or preserved [68, 46, 10, 66]. Some others defined more complex criteria such as saliency and sensitivity [32, 34, 72].
Once a pruning criterion is established, we can assess whether a model is easy or difficult to prune based on that criterion. For instance, using a simple $l_{2}$ norm criterion, a model whose parameters exhibit either very large or very small $l_{2}$ norms is clearly easier to prune compared to one with uniformly distributed $l_{2}$ norms. From this viewpoint, we can classify models as more or less suitable for pruning. This concept forms the foundation of our work.
Previously, researchers have generally approached pruning improvement in two ways: either by refining the pruning criteria or by transforming the model from “bad” to “good” according to a given criterion. The first approach is more straightforward and has been extensively studied in prior work. In contrast, the second approach has received comparatively less attention. However, several studies adopting this latter perspective have achieved promising results. Notably, sparsity regularization methods [61, 9, 60] exemplify this approach by encouraging models to become sparser, thereby facilitating more effective pruning.
Our work follows the second approach and introduces a novel method inspired by sparsity regularization. We focus on this approach because, despite the many pruning criteria proposed, most lack solid theoretical foundations and have become overly complex. Therefore, in this work, we adopt a simple variant of the $l_{2}$ norm—the group norm from Fang et al. [9]—as our pruning criterion. Additionally, we bring the concept of metanetworks from meta-learning into pruning. A metanetwork is a network that takes another neural network as input and produces a modefied neural network as output. In our method, the metanetwork transforms a “bad” network into a “good” one, making it more amenable to pruning. Once trained, our metanetwork-based approach requires only a feedforward pass through the metanetwork and some standard finetuning to achieve state-of-the-art pruning results.
Our contributions can be summarized as follows:
  • Introducing the idea of metanetwork in meta-learning into pruning for the first time.
  • Proposing a novel pruning method that is simple and powerful.
  • Achieving state-of-the-art performance on several popular pruning tasks.
It is worth noting that this paper focuses solely on structural pruning, which does not rely on specialized AI accelerators or software to reduce memory usage and computational costs. This makes it more versatile and broadly applicable. However, all our methods can also be applied directly to unstructured pruning.
2 Related Work
2.1 Metanetworks
Using neural networks to process other neural networks has emerged as an intriguing research direction. Early studies demonstrated that neural networks can extract useful information directly from the weights of other networks [58]. We define a metanetwork as a neural network that takes another neural network as input and outputs either information about it or a modified network. Initial metanetworks were simple multilayer perceptrons (MLPs), while more recent designs incorporate stronger inductive biases by preserving symmetries [13] inherent in the input networks. Broadly, metanetwork architectures can be categorized into two perspectives. The weight space view applies specially designed MLPs directly on the model’s weights [71, 48, 73, 57, 74, 75, 56]. In contrast, the graph view transforms the input network into a graph and applies graph neural networks (GNNs) to it [40, 30, 27]. In this work, we adopt the graph view to build our metanetwork, leveraging the natural correspondence between graph nodes and neurons in a network layer, as well as the symmetries they exhibit [47]. Given the maturity and effectiveness of GNNs, the graph metanetwork approach offers both elegance and practicality.
2.2 Graph Neural Networks
Graph neural networks (GNNs) are a class of neural networks specifically designed to operate on graph-structured data, where graphs consist of nodes and edges with associated features [63, 52, 29]. GNNs have gained significant attention in recent years, with frameworks such as the message passing neural network (MPNN) [12] proving to be highly effective. In this paper, we employ the message passing framework, specifically using Principal Neighborhood Aggregation (PNA) [3] as the backbone architecture for our metanetwork.
2.3 Sparsity Regularization-Based Pruning
Networks that are inherently sparse tend to be easier to prune effectively. Prior research has explored various regularization techniques to encourage sparsity in neural networks to facilitate pruning. This idea of sparsity regularization has been widely adopted across many works [61, 14, 25, 44, 22, 43, 64, 65, 9, 60]. Our method inherits this principle to some extent, as our metanetwork can be viewed as transforming the original network into a sparser version that is more amenable to pruning.
2.4 Learning to Prune & Meta Pruning
Given the complexity of pruning, several previous works have leveraged neural networks [5, 22, 62, 1], reinforcement learning [50, 70, 23], and other techniques to automatically learn pruning strategies. Viewing pruning as a learning problem naturally leads to the idea of meta-learning, which learns “how to learn” [24]. Some prior works have specifically applied meta-learning to pruning [45, 37]. While our work also adopts a meta-learning framework, it differs fundamentally from these previous approaches by using a metanetwork.
3 Method
Our approach centers on training a metanetwork that learns the pruning strategy automatically and can transform a neural network that is difficult to prune into one that is easier to prune, thereby improving pruning effectiveness and efficiency.
We begin by describing how our metanetwork transforms one network into another, including the conversion between network and graph, and the design of our metanetwork(a graph neural network). Next, we introduce the group $l_{2}$ norm criterion for meta-training and pruning. Then we detail the meta-training pipeline used to train the metanetwork. Finally, we present the complete pruning process.
[…]
3.1 Equivalent Conversion between Neural Network and Graph
To enable a neural network to be processed by another network, we first need to represent the original network in a structured data format. In this work, we convert the neural network into a graph, drawing inspiration from prior works such as Kofinas et al. [30], Lim et al. [40], but with notable differences in implementation.
To convert a network into a graph, we establish the correspondence between the components of a neural network and elements of a graph as follows:
  • Node: Each neuron or convolutional channel in the network is represented as a node in the graph.
  • Edge: An edge exists between two nodes if there is a direct connection between them in the original network. This includes connections between neurons or channels within linear or convolutional layers, as well as residual connections.
  • Node Features: Node features comprise parameters associated with neurons, here we use the weight, bias, running mean, running var of batchnorm [26], including both batchnorms from previous adjacent layers and previous skip connection layers (if exist).
  • Edge Features: Edge features encode the weights of connections between nodes. We treat linear connections and residual connections as special cases of convolutional connections with a kernel size of 1. For a $k\times k$ convolution between two channels, we linearize it into a $k^{2}$-dim edge feature vector.
Since graph neural networks require all edge features to have the same dimension, we adopt one of two strategies to standardize them:
  • Pad all convolutional kernels to the same size (e.g., the maximum kernel size in the network), then flatten them to form fixed-length edge feature vectors.
  • Flatten convolutional kernels of varying sizes, then apply learned linear transformations to project them into a uniform feature dimension.
By following these rules, we can generate a graph that corresponds to a given neural network. Conversely, the inverse transformation can be applied to convert a graph back into an equivalent network. In essence, we establish an equivalent conversion between networks and graphs. For a more intuitive understanding, see Figure 1.
Figure 1: We visualize a two layer small CNN as example. Our graph contains 6 nodes corresponding with 6 neurons in the network. For neurons connected via convolutional channel like (1,3) and connected by residual connection like (1,5), we add an edge between their corresponding nodes. Then we generate node and edge features as shown in the figure. For some special cases, we use default values to replace non-existent values. For example, for node 3 who doesn’t have a residual connection, we set the last 4 number of its node feature as [1, 0, 0, 1]. This means we treat it as with previous residual batchnorm of weight 1, bias 0, running mean 0, running var 1, performing the same as with no batchnorm.
3.2 Graph metanetwork design
After changing the input network into a graph, we use our metanetwork, a graph neural network, to process the input graph. We build our metanetwork based on the message passing framework [12] and PNA [3] architecture.
We use $v_{i}$ to represent node $i$’s feature vector, and $e_{ij}$ to represent feature vector of the edge between $i$ and $j$. We consider undirected edges, i.e., $e_{ij}$ is the same as $e_{ji}$. Our metanetwork is as follows.
First, all input node and edge features are encoded into the same hidden dimension.
$v\leftarrow\text{MLP}_{NodeEnc}(v^{in}),$ (1)
$e\leftarrow\text{MLP}_{EdgeEnc}(e^{in}).$ (2)
Then they will pass through several message passing layers. In each layer, we generate the messages:
$m_{ij}$ $\leftarrow\text{MLP}_{Node}^{1}(v_{i})\odot\text{MLP}_{Node}^{2}(v_{j})\odot e_{ij},$ (3)
$m^{\prime}_{ij}$ $\leftarrow\text{MLP}_{Node}^{1}(v_{j})\odot\text{MLP}_{Node}^{2}(v_{i})\odot(e_{ij}\odot EdgeInvertor),$ (4)
where $m_{ij}$ and $m^{\prime}_{ij}$ respectively encode message from $i$ to $j$ and message from $j$ to $i$, and EdgeInvertor is defined as:
$EdgeInvertor\triangleq[\underbrace{1,1,\dots,1}_{\text{hidden\_dim}/2},\underbrace{-1,-1,\dots,-1}_{\text{hidden\_dim}/2}].$ (5)
The intuition behind EdgeInvertor is to let the first half dimensions to learn undirectional features invariant to exchanging $i,j$, and the second half to capture directional information that changes equivariantly with reverting $i,j$. Empirically we found this design the most effective among various ways to encode edge features.
Then, we aggregate the messages to update node features:
$v_{i}\leftarrow v_{i}+\text{PNA}_{Aggr}(m_{ij})+\text{PNA}_{Aggr}(m_{ij}^{\prime}),$ (6)
where we have
$\text{PNA}_{Aggr}(m_{ij})\triangleq\text{MLP}_{Aggr}\left(\left[\mathop{\text{MEAN}}\limits_{j:(i,j)\in E}(m_{ij}),\mathop{\text{STD}}\limits_{j:(i,j)\in E}(m_{ij}),\mathop{\text{MAX}}\limits_{j:(i,j)\in E}(m_{ij}),\mathop{\text{MIN}}\limits_{j:(i,j)\in E}(m_{ij})\right]\right).$ (7)
For edge features in each layer, we also update them by:
$e_{ij}\leftarrow e_{ij}$ $+\text{MLP}_{Edge}^{1}(v_{i})\odot\text{MLP}_{Edge}^{2}(v_{j})\odot e_{ij}$ (8)
$+\text{MLP}_{Edge}^{1}(v_{j})\odot\text{MLP}_{Edge}^{2}(v_{i})\odot(e_{ij}\odot EdgeInvertor).$
Finally, we use a decoder to recover node and edge features to their original dimensions:
$v_{pred}\leftarrow\text{MLP}_{NodeDec}(v),$ (9)
$e_{pred}\leftarrow\text{MLP}_{EdgeDec}(e).$ (10)
We multiply our predictions by a residual coefficient and add them to the original inputs as our outputs:
$v_{out}=\alpha\cdot v_{pred}+v_{in},$ (11)
$e_{out}=\beta\cdot e_{pred}+e_{in},$ (12)
where $\alpha$ and $\beta$ in practice are set to a small real numbers like $0.01$. This design enables the metanetwork to learn only the delta weights on top of the original weights, instead of a complete reweighting.
3.3 Group $l_{2}$ Norm Criterion
We use group $l_{2}$ norm from Fang et al. [9] as our pruning criterion. The design philosophy of group $l_{2}$ norm is to measure the sparsity of pruning groups. In structural pruning, we only prune a group of elements in a network. For example, if we prune a neuron, then all edges connected to the neuron should also be pruned, which we say that they are in a group. A more detailed explanation of group can be found in Fang et al. [9]. Since structural pruning is performed on groups, traditional $l_{2}$ norm that treats nodes and edges in isolation is no longer viable. Group $l_{2}$ norm that takes a whole group into consideration has shown to be more effective.
Given a parameter group $g=\{w_{1},w_{2},\ldots,w_{|g|}\}$. For each parameters $w$ with $K$ prunable dimensions indexed by $w[k]$, we measure the importance of the k-th prunable dimension by :
$I_{g,k}=\sum_{w\in g}\|w[k]\|^{2}$ (13)
During training, our sparsity loss is defined as :
$\mathcal{R}(g,k)=\sum_{k=1}^{K}\gamma_{k}\cdot I_{g,k}=\sum_{k=1}^{K}\sum_{w\in g}\gamma_{k}\|w[k]\|^{2}$ (14)
Where $\gamma_{k}$ refers to a shrinkage strength applied to the parameters to modify the gradients for better training. Defined as :
$\gamma_{k}=2^{\alpha(I_{g}^{\max}-I_{g,k})(I_{g}^{\max}-I_{g}^{\min})}$ (15)
And we use $\alpha=4$ in all our experiments.
When pruning, we calculate the normalized relative score of all prunable dimensions in all groups
$\hat{I}_{g,k}=N\cdot I_{g,k}/\sum\{\text{TopN}(I_{g})\}$ (16)
where $N$ is a hyperparameter that chooses the top $N$ importance scores in the group $g$ to normalize every prunable dimension $k$’s score $I_{g,k}$. We prune unimportant dimensions of groups globally from lower to higher according to the normalized relative score.
Figure 2: Meta-Training Pipeline.
3.4 Meta-Training Pipeline
The goal of training our metanetwork to enable it to transform a hard-to-prune network into an easy-to-prune one, referred to as meta-training. This section details our meta-training pipeline, which is illustrated in Figure 2.
To prepare for meta-training, first we need the training data. Meta-training data is comprised of two parts. The first is a set of neural networks to prune, typically between 2 to 8, which serve as the origin network in meta-training. The second is the traditional training datasets, such as CIFAR-10, CIFAR-100, or ImageNet, used to calculate the accuracy loss.
During each training iteration, we select one model from data models as the origin network. The origin network is then converted into its graph representation. This graph is passed through the metanetwork, which outputs a new graph. We subsequently convert this new graph back into a new network.
Two types of losses are computed on the new network:
1) Accuracy Loss: We feed the training data into the new network and compute the cross-entropy loss based on the output predictions, ensuring the new network to have high accuracy after finetuning.
2) Sparsity Loss: We apply the group $l_{2}$ norm (as defined in 14) on the new network to encourage structured sparsity, making it more amenable to pruning.
The gradients from the loss are backpropagated through the metanetwork to update its parameters.
The accuracy loss ensures that the metanetwork preserves the performance of the original network, while the sparsity loss encourages the metanetwork to generate a network with group sparsity that facilitates pruning. After meta-training, our metanetwork can generate a transformed network that has almost no loss in accuracy after finetuning, and is significantly easier to prune.
Empirically, we observe that the metanetwork performs much better on hard-to-prune origin networks. However, most networks are initially easy to prune—for example, ResNet-56 on CIFAR-10 can be pruned with a 1.3× speedup and finetuned without accuracy loss. Therefore, in all our experiments (both during meta-training and pruning), we first prune the network slightly—without any accuracy degradation—to make it harder to prune. This pre-pruning step allows us to fully leverage the capability of the metanetwork.
3.5 Pruning a Target Network
Our pruning pipeline only involves a feedforward through metanetwork and finetuning. It can be summarized as:
$InitialPruning(Finetuning)\rightarrow Metanetwork(Finetuning)\rightarrow Pruning(Finetuning)$
As discussed in 3.4, we first do initial pruning to make our network harder to prune while maintaining the accuracy. Then we feed the hard-to-prune network through the metanetwork to transform it into a much easier-to-prune one, which also has almost no accuracy loss after finetuning. In the end, we perform the final pruning to reduce its size, and the final finetuning to evaluate the performance.
4 Experiments
4.1 Setup
We carry out our experiments on three most classical, popular, and representative image recognition tasks, including resent56 [16] on CIFAR10 [31], VGG19 [54] on CIFAR100 [31] and resnet50 [16] on ImageNet [4]. See appendix A for more general setups.
[…]
4.2 ResNet56 on CIFAR10
Table 1: ResNet56 on CIFAR10 (see full table in appendix 4)
Method Base Pruned $\Delta$ Acc Pruned FLOPs Speed Up
FPGM [19] 93.59 92.93 -0.66 52.6% 2.11
FPC [21] 93.59 93.24 -0.35 52.9% 2.12
DMC [11] 93.62 92.69 -0.93 50.0% 2.00
GNN-RL [70] 93.49 93.59 0.10 54.0% 2.17
DepGraph w/o SL [9] 93.53 93.46 -0.07 52.6% 2.11
DepGraph with SL [9] 93.53 93.77 0.24 52.6% 2.11
ATO [62] 93.50 93.74 0.24 55.0% 2.22
Meta-Pruning (ours) 93.51 94.08 0.57 56.5% 2.30
C-SGD [7] 93.39 93.44 0.05 60.8% 2.55
Greg-1 [60] 93.36 93.18 -0.18 60.8% 2.55
Greg-2 [60] 93.36 93.36 0.00 60.8% 2.55
DepGraph w/o SL [9] 93.53 93.36 -0.17 60.2% 2.51
DepGraph with SL [9] 93.53 93.64 0.11 61.1% 2.57
ATO [62] 93.50 93.48 -0.02 65.3% 2.88
Meta-pruning (ours) 93.51 93.69 0.18 65.6% 2.91
Meta-pruning (ours) 93.51 93.25 -0.26 66.9% 3.02
Table 1 shows our results of pruning ResNet56 [16] on CIFAR10 [31]. Our algorithm outperforms all previous works. We can prune 65.6% FLOPs (i.e. 2.91x Speed Up) without any loss of accuracy, which is quite amazing. See Appendix B for more results and implementation details.
4.3 VGG19 on CIFAR100
Table 2: VGG19 on CIFAR100 (see full table in appendix 7)
Method Base Pruned $\Delta$ Acc Pruned FLOPs Speed Up
OBD [59] 73.34 60.66 -12.68 83.58% 6.09
EigenD [59] 73.34 65.18 -8.16 88.64% 8.80
Greg-2 [60] 74.02 67.75 -6.27 88.69% 8.84
DepGraph w/o SL [9] 73.50 67.60 -5.44 88.73% 8.87
DepGraph with SL [9] 73.50 70.39 -3.11 88.79% 8.92
Meta-Pruning (ours) 73.65 70.06 -3.59 88.85% 8.97
Table 2 shows our results of pruning VGG19 [54] on CIFAR100 [31]. Our algorithm reaches 70.06% pruned accuracy with 88.85% FLOPs pruned (i.e. 8.97x Speed Up). While our pruning speed up is larger than all previous works, our pruned accuracy is only a little bit lower than DepGraph with SL and higher than all other works. See Appendix C for more results and implementation details.
4.4 ResNet50 on ImageNet
Table 3: ResNet50 on ImageNet (see full table in appendix 10)
Method Base Top-1(Top-5) Pruned Top-1($\Delta$) Pruned Top-5($\Delta$) Pruned FLOPs
DCP [77] 76.01%(92.93%) 74.95%(-1.06%) 92.32%(-0.61%) 55.6%
CCP [49] 76.15%(92.87%) 75.21%(-0.94%) 92.42%(-0.45%) 54.1%
FPGM [19] 76.15%(92.87%) 74.83%(-1.32%) 92.32%(-0.55%) 53.5%
ABCP [42] 76.01%(92.96%) 73.86%(-2.15%) 91.69%(-1.27%) 54.3%
DMC [11] 76.15%(92.87%) 75.35%(-0.80%) 92.49%(-0.38%) 55.0%
Random [38] 76.15%(92.87%) 75.13%(-1.02%) 92.52%(-0.35%) 51.0%
DepGraph [9] 76.15%(-) 75.83%(-0.32%) - 51.7%
ATO [62] 76.13%(92.86%) 76.59%(+0.46%) 93.24%(+0.38%) 55.2%
DTP [39] 76.13%(-) 75.55%(-0.58%) - 56.7%
ours 76.14%(93.11%) 76.17%(+0.03%) 92.94%(-0.17%) 56.9%
Table 3 shows our results of pruning ResNet50 [16] on ImageNet[4]. Our algorithm reaches 76.17% top-1 and 92,95% top-2 accuracy with 56.9% FLOPs pruned. With the largest pruned FLOPs, our pruned accuracy is higher than almost all previous works (including DTP [39]). The only exception is that our accuracy is a bit lower than ATO [62], but our pruned FLOPs is much larger (1.7%) than it. See appendix D for more results and implementation details.
4.5 Ablation Study
How metanetwork works. To help readers visually grasp the effect of metanetwork, we visualized the "Test Acc VS. Speed Up" curve of both origin network and network after feedforward through metanetwork and finetuning . We generate the curve by prune the network little by little. After each pruning step, we evaluated the model’s performance on the test set (no finetuning in this step) and recorded a data point representing the current speed up and corresponding test accuracy. Connecting these points forms a curve that illustrates how accuracy changes as the speed up increases. The "Test Acc VS. Speed Up" curve of ResNet on CIFAR10 can be seen in figure 3(a). Similar curves of VGG19 on CIFAR100 and ResNet50 on ImageNet can be seen in appendix E.1.
As shown in the "Test Accuracy vs. Speed Up" curves, as the pruning speed up increases, the accuracy drops at a significantly slower rate after applying the metanetwork and finetuning. Take figure 3(a) as an example (this is exactly the metanetwork we use in 4.2). For the original network, the accuracy decreases rapidly and nearly collapses before reaching a speed up of 1.50x, suggesting that directly using group l2 norm criterion 3.3 to prune the origin network removes many important parts of the network (i.e. the origin network is hard to prune under this criterion). However, after metanetwork and finetuning, pruning to 2.0x speed up without finetuning causes almost no drop in accuracy, suggesting that using group l2 norm criterion on this network can efficiently distinguish the unimportant parts of this network (i.e. the network after metanetwork and finetuning is much easier to prune under this criterion). This supports our claim in 3 that metanetwork can transform a neural network that is difficult to prune into one that is much easier to prune.
(a) Origin VS After Metanetwork
(b) Impact Of Training Epochs On Metanetworks
Figure 3: Visualization of how metanetworks effect the networks (ResNet56 on CIFAR10)
Trade-off between accuracy and speed up. We observe an interesting phenomenon during meta-training: as the number of training epochs increases, the metanetwork’s ability to make the target network easier to prune becomes stronger. However, this comes at the cost of its ability to maintain the accuracy. As shown in figure 3(b), with more training epochs, the "Test Accuracy vs. Speed Up" curve shifts downward, and the flat portion of the curve becomes longer. This characteristic allows us to adaptively meet different pruning requirements. If a higher level of pruning is desired and a moderate drop in accuracy is acceptable, a metanetwork trained for more epochs would be preferable. Conversely, if preserving accuracy is more important, a metanetwork trained for fewer epochs would be a better choice.
Finetuning after Metanetwork. See appendix E.2 for our study on the impact of the number of finetuning epochs after metanetwork on performance.
5 Conclusion
In this paper we proposes a brand new meta-learning framework for network pruning. For the first time we introduce the concept of metanetwork (a network that takes another network as input and produces a modified network as output) into pruning. The metanetwork is trained to automatically learn effective pruning strategies. After training, it can take a network that is hard to prune as input and generate a transformed network that is much easier to prune. During pruning, our method needs nothing more than a feedforward through metanetwork and trivial finetuning to prune at state-of-the-art.
Advantages Over Previous Works. Our metanetwork is capable of considering the weights of multiple nearby layers, which sets it apart from most previous approaches. Unlike earlier methods that either focus only on local weights or incorporate more layers but rely on simple, hand-crafted relationships between them, our metanetwork automatically learns complex relationships among neighboring layers. These relationships are too intricate to be expressed by simple human-defined rules. So we use metanetwork to learn it automatically during training. This can explain why our work outperforms almost all previous works,
Limitations And Future Work. Our work also has several limitations. First, converting networks into graph representations requires GPUs with high VRAM, making it challenging to apply our method to large and densely parameterized models. Second, our current approach can’t be directly applied to networks with different architectures, such as Transformers or RNNs without appropriate modifications. However, these limitations also present opportunities for future research. We think the following directions are promising and worth exploring: (1) Extending our method to accommodate diverse architectures, including Transformers, diffusion models, and others. (2) Investigating alternative approaches to constructing the metanetworks. (3) Applying similar meta-learning frameworks to a broader range of tasks.
Acknowledgments and Disclosure of Funding
Use unnumbered first level headings for the acknowledgments. All acknowledgments go at the end of the paper before the list of references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found at: https://neurips.cc/Conferences/2025/PaperInformation/FundingDisclosure.
Do not include this section in the anonymized submission, only in the final paper. You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
Appendix A General Experiment Setup
A.1 General Settings
Optimizer: For all data model training and fine-tuning, we use torch.optim.SGD with momentum=0.9. For meta-training, we employ torch.optim.AdamW. All learning rates are controlled using torch.optim.lr_scheduler.MultiStepLR with gamma=0.1. The learning rate is adjusted at specified milestones; for instance, if milestones=[50, 90], the learning rate is multiplied by 0.1 at epochs 50 and 90. These settings are not strictly necessary—other optimizers and schedulers may be equally effective.
Data for Meta-Training: As discussed in Section 3.4, meta-training requires two types of data: data models and traditional training datasets. To generate a data model, we train a base model, perform initial pruning followed by finetuning, and finally save the model along with relevant information. Typically, generating 3–10 data models is sufficient, with 2–8 used for meta-training and the remainder reserved for meta-evaluation or testing.
Meta Evaluation (Meta Eval): Meta evaluation is a process used to assess the quality of the metanetwork during meta-training. We did not include it in our final experiments due to its high computational cost; however, we retain it as an optional tool for future research. At the end of each meta-training epoch, we evaluate the metanetwork by feeding unseen data models (not used in meta-training) through the network, finetuning, and then pruning to achieve the maximum speed-up while maintaining accuracy above a predefined threshold. A higher resulting speed-up indicates better metanetwork performance.
Visualizing Metanetworks: This is a key technique used throughout our experiments. When we refer to visualizing a metanetwork, we mean passing a model through the metanetwork, followed by finetuning, and then visualize the “Test Accuracy vs. Speed-Up” curve of the resulting model, as illustrated in Figure 3. This visualization provides insights into the metanetwork’s behavior and helps determine whether it is suitable for pruning. An ideal metanetwork should exhibit a long flat region in the curve where accuracy remains close to or above the target pruning accuracy. There are two ways to select a suitable metanetwork: (1) using meta-evaluation during meta-training, or (2) visualizing metanetworks post-training. In our final experiments, we exclusively use the second method, which is significantly more efficient. We typically do not visualize all metanetworks but instead search for the best one using a binary search strategy.
Big Batch Size and Small Batch Size: We use two different batch sizes in our experiments: a small batch size and a big batch size. The big batch size is used during meta-training to compute accuracy loss on new networks. Due to memory constraints, such large batches cannot be processed in a single forward and backward pass. Instead, we accumulate gradients over multiple smaller mini-batches before performing a single optimizer update—a common PyTorch practice. The big batch size is thus expressed as batchsize * iters, indicating that one optimizer.step() is performed after iters forward/backward passes, each using a mini-batch of size batchsize. For all other standard training and finetuning tasks, we use the small batch size.
A.2 General Meta Training Details
Equivalent Conversion between Network and Graph: During meta-training, we first convert the original network into an origin graph. Then, we feed this origin graph through the metanetwork to generate a new graph. Finally, we convert the new graph back into a new network. When converting the original network into the origin graph, all network parameters are mapped to either node or edge features in the graph. However, some graph features do not correspond to any network parameter. For these missing features, we use default values that effectively simulate the absence of those parameters. When converting the new graph back into the new network, we only map those trainable parameters from the new graph back to the new network. For untrainable parameters in the new network, we keep it the same as the old network. We provide more explanations for this conversion process in the following subsections: B, C, and D.
The Relative Importance of Sparsity Loss and Accuracy Loss: In our experiments, we introduce a hyperparameter called "pruner reg", which controls the relative importance of sparsity loss compared to accuracy loss. During backpropagation, the gradient from the sparsity loss is scaled(multiplied) by this "pruner reg" value, while the gradient from the accuracy loss remains unscaled.
Meta Training Milestone: As discussed in the Ablation Study 3, as the number of training epochs increases, the metanetwork’s ability to make the network easier to prune becomes stronger, while its ability to maintain the accuracy becomes weaker. During meta-training, we usually set a milestone for our learning rate scheduler. Before this milestone, we use a relatively large learning rate to quickly improve the metanetwork’s ability to make networks more prunable. After the milestone, we switch to a smaller learning rate to finetune the metanetwork and slow down the change of the two abilities, allowing us to select a well-balanced metanetwork. To determine a reasonable milestone value, we can first perform a short meta-training phase using the large learning rate, then visualize the resulting metanetworks to identify the point where the accuracy-maintaining ability of metanetwork is a bit stronger than we expect.
Choose the Appropriate Metanetwork: During meta-training, we save the metanetwork after each epoch. After training is complete, we search for the most suitable metanetwork by visualizing its performance using a binary search strategy. Specifically, we start by visualizing the metanetwork from the middle epoch. If the accuracy of its flat region is below our target pruned accuracy, we next visualize the metanetwork from the first quarter of training. Otherwise, we check the one from the third quarter. We continue this process iteratively until we find a metanetwork that meets our criteria: a relatively high accuracy in the flat region and a long flat region indicating robustness to pruning.
A.3 General Pruning Details :
Pruning Speed Up: Once a suitable metanetwork has been selected for pruning, the next step is to determine the target speed up. In general, after visualizing the metanetwork, we observe a flat region followed by a sharp decline in accuracy. We define a turning point as the speed up at which the accuracy drops below a certain threshold. Experimentally, we find that we can safely prune the model at a speed up slightly higher than this turning point without incurring any accuracy loss after fine-tuning. For instance, when pruning ResNet-56 on CIFAR-10, we consider an accuracy drop below 0.93 as the turning point (typically around 2.0$\times$ speed up), and in practice, we are able to achieve a 2.9$\times$ speed-up with no loss in accuracy after finetuning.
Appendix B ResNet56 on CIFAR10
B.1 Equivalent Conversion between Network and Graph
We change ResNet56 into a graph with node featrues of 8 dimensions and edge features of 9 dimensions.
The node features consist of 4 features derived from the batch normalization parameters (weight, bias, running mean, and running variance) of the previous layer, along with 4 features from the batch normalization of the previous residual connection. If no such residual connection exists, we assign default values $[1,0,0,1]$ to the corresponding 4 features.
The edge features are constructed by zero-padding all convolutional kernels to a size of $3\times 3$ (note that our networks only contain kernels of size $3\times 3$ or $1\times 1$), and then flattening them into feature vectors. We treat all residual connections as convolutional layers with kernel size $1\times 1$. For example, consider two layers A and B connected via a residual connection, with neurons indexed as $1,2,3,\ldots$ in both layers. If the residual connection has no learnable parameters (i.e., it directly adds the input to the output), we represent the edge from $A_{i}$ to $B_{i}$ as a $1\times 1$ convolutional kernel with value 1, and the edge from $A_{i}$ to $B_{j}$ (where $i\neq j$) as a $1\times 1$ kernel with value 0. If the residual connection includes parameters (e.g., a downsample with a convolutional layer and batch normalization), we construct the edge features in the same way as for standard convolutional layers.
B.2 Meta Training
Data Models : We generate 10 models as our data models, among them, 8 are used for meta-training and 2 are used for validation (visualize the metanetwork). When generating each data model, we train them 200 epochs with learning rate 0.1 and milestone "100, 150, 180", then pruning with speed up 1.32x, followed by a finetuning for 80 epochs with learning rate 0.01 and milestone "40, 70". Finally, we can get a network with accuracy around 93.5%.
One Meta Training Epoch : Every epoch we enumerate over all 8 data models. When enumerate a data model, we feedforward it through the metanetwork and generate a new network. We feed all our CIFAR10 training data into the new network to calculate the accuracy loss, and use the parameters of new network to calculate the sparsity loss. Then we backward the gradients from both two losses to update our metanetwork.
Training: We train our metanetwork with learning rate 0.001, milestone "12", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 28 as our final network for pruning.
B.3 Full Results (Including Error Bars)
All results are summarized in Table 4, where we repeat the pruning process 3 times using different seeds: 7, 8, 9. It is important to note that due to the design of our algorithm and implementation, we cannot prune the network to exactly the same speed up across different runs. For instance, when targeting a 1.3$\times$ speed up, the actual achieved speed up may be slightly larger, such as 1.31$\times$, 1.32$\times$, or 1.35$\times$.
The aggregated statistical results are presented in Table 5, where each value is reported in the form of mean(standard deviation).
Table 4: ResNet56 on CIFAR10 Full
Method Base Pruned $\Delta$ Acc Pruned FLOPs Speed Up
NISP [69] 43.2% 1.76
Geometric [20] 93.59 93.26 -0.33 41.2% 1.70
Polar [76] 93.80 93.83 0.03 46.8% 1.88
DCP-Adapt [77] 93.80 93.81 0.01 47.0% 1.89
CP [35] 92.80 91.80 -1.00 50.0% 2.00
AMC [23] 92.80 91.90 -0.90 50.0% 2.00
HRank [41] 93.26 92.17 -1.09 50.0% 2.00
SFP [18] 93.59 93.36 -0.23 52.6% 2.11
ResRep [8] 93.71 93.71 0.00 52.8% 2.12
SCP [28] 93.69 93.23 -0.46 51.5% 2.06
FPGM [19] 93.59 92.93 -0.66 52.6% 2.11
FPC [21] 93.59 93.24 -0.35 52.9% 2.12
DMC [11] 93.62 92.69 -0.93 50.0% 2.00
GNN-RL [70] 93.49 93.59 0.10 54.0% 2.17
DepGraph w/o SL [9] 93.53 93.46 -0.07 52.6% 2.11
DepGraph with SL [9] 93.53 93.77 0.24 52.6% 2.11
ATO [62] 93.50 93.74 0.24 55.0% 2.22
Meta-Pruning (ours) 93.51 94.08 0.57 56.5% 2.30
Meta-Pruning (ours) 93.51 93.56 0.05 56.5% 2.30
Meta-Pruning (ours) 93.51 93.81 0.30 56.9% 2.32
GBN [67] 93.10 92.77 -0.33 60.2% 2.51
AFP [6] 93.93 92.94 -0.99 60.9% 2.56
C-SGD [7] 93.39 93.44 0.05 60.8% 2.55
Greg-1 [60] 93.36 93.18 -0.18 60.8% 2.55
Greg-2 [60] 93.36 93.36 0.00 60.8% 2.55
DepGraph w/o SL [9] 93.53 93.36 -0.17 60.2% 2.51
DepGraph with SL [9] 93.53 93.64 0.11 61.1% 2.57
ATO [62] 93.50 93.48 -0.02 65.3% 2.88
Meta-pruning (ours) 93.51 93.69 0.18 65.6% 2.91
Meta-pruning (ours) 93.51 93.42 -0.09 66.0% 2.94
Meta-pruning (ours) 93.51 93.31 -0.20 66.6% 2.99
Meta-pruning (ours) 93.51 93.25 -0.26 66.9% 3.02
Meta-pruning (ours) 93.51 93.20 -0.31 66.7% 3.00
Meta-pruning (ours) 93.51 93.18 -0.33 67.5% 3.08
Table 5: ResNet56 on CIFAR10 Statistics
Pruned Accuracy Pruned FLOPs Speed Up
93.82($\pm$0.21) 56.63%($\pm$0.19%) 2.3067($\pm$0.0094)
93.47($\pm$0.16) 66.07%($\pm$0.41%) 2.9467($\pm$0.0330)
93.21($\pm$0.03) 67.03%($\pm$0.34%) 3.0333($\pm$0.0340)
B.4 Computer Resources and Hyperparameters
See Table 6.
Table 6: ResNet56 on CIFAR10 Hyperparameters
Types Name Value
Compute Resources GPU NVIDIA RTX 4090
parallel No
Batch size small batch size 128
big batch size 500 $\times$ 100
Prepare Data Models data model num 10 ( 8 + 2 )
Train From Scratch epoch 200
lr 0.1
weight decay 0.0005
milestone "100, 150, 180"
Initial Pruning speed up 1.32
Finetuning epoch 80
lr 0.01
weight decay 0.0005
milestone "40, 70"
Metanetwork num layer 8
hiddim 64
in node dim 8
in edge dim 9
node res ratio 0.01
edge res ratio 0.01
Meta Training lr 0.001
weight decay 0.0005
milestone "12"
pruner reg 10
Final Pruning metanetwork epoch 28
speed up 2.3, 2.9, 3.0
Finetuning After Metanetwork epoch 100
lr 0.01
weight decay 0.0005
milestone "60, 90"
Finetuning After Pruning epoch 140
lr 0.01
weight decay 0.0005
milestone "80, 120"
Appendix C VGG19 on CIFAR100
C.1 Equivalent Conversion between Network and Graph
We change VGG19 into a graph with node featrues of 5 dimensions and edge features of 9 dimensions.
The node features consist of 4 features derived from the batch normalization parameters (weight, bias, running mean, and running variance) of the previous layer and 1 feature derived from the bias of previous layer (0 if bias doesn’t exist).
The edge features are constructed by zero-padding all convolutional kernels to a size of $3\times 3$ (note that our networks only contain kernels of size $3\times 3$ or $1\times 1$), and then flattening them into feature vectors.
C.2 Meta Training
Data Models : We generate 10 models as our data models, among them, 8 are used for meta-training and 2 are used for validation (visualize the metanetwork). When generating each data model, we train them 200 epochs with learning rate 0.1 and milestone "100, 150, 180", then pruning with speed up 2.0x, followed by a finetuning for 140 epochs with learning rate 0.01 and milestone "80, 120". Finally, we can get a network with accuracy around 73.5%.
One Meta Training Epoch : Every epoch we enumerate over all 8 data models. When enumerate a data model, we feedforward it through the metanetwork and generate a new network. We feed all our CIFAR100 training data into the new network to calculate the accuracy loss, and use the parameters of new network to calculate the sparsity loss. Then we backward the gradients from both two losses to update our metanetwork.
Training: We train our metanetwork with learning rate 0.001, milestone "15", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 51 as our final network for pruning.
C.3 Full Results (Including Error Bars)
All results are summarized in Table 7, where we repeat the pruning process 3 times using different seeds: 7, 8, 9. It is important to note that due to the design of our algorithm and implementation, we cannot prune the network to exactly the same speed up across different runs. For instance, when targeting a 8.95$\times$ speed up, the actual achieved speed up may be slightly larger, such as 8.95$\times$, 9.01$\times$, or 9.02$\times$.
The aggregated statistical results are presented in Table 8, where each value is reported in the form of mean(standard deviation).
Table 7: VGG19 on CIFAR100 Full
Method Base Pruned $\Delta$ Acc Pruned FLOPs Speed Up
OBD [59] 73.34 60.70 -12.64 82.55% 5.73
OBD [59] 73.34 60.66 -12.68 83.58% 6.09
EigenD [59] 73.34 65.18 -8.16 88.64% 8.80
Greg-1 [60] 74.02 67.55 -6.67 88.69% 8.84
Greg-2 [60] 74.02 67.75 -6.27 88.69% 8.84
DepGraph w/o SL [9] 73.50 67.60 -5.44 88.73% 8.87
DepGraph with SL [9] 73.50 70.39 -3.11 88.79% 8.92
Meta-Pruning (ours) 73.65 70.06 -3.59 88.85% 8.97
Meta-Pruning (ours) 73.65 69.19 -4.46 88.91% 9.02
Meta-Pruning (ours) 73.65 67.05 -6.60 89.07% 9.15
Table 8: VGG19 on CIFAR100 Statistics
Pruned Accuracy Pruned FLOPs Speed Up
68.77($\pm$1.26) 88.94%($\pm$0.09%) 9.0467($\pm$0.0759)
C.4 Computer Resources and Hyperparameters
See Table 9.
Table 9: VGG19 on CIFAR100 Hyperparameters
Types Name Value
Compute Resources GPU NVIDIA RTX 4090
parallel No
Batch size small batch size 128
big batch size 500 $\times$ 100
Prepare Data Models data model num 10 ( 8 + 2 )
Train From Scratch epoch 200
lr 0.1
weight decay 0.0005
milestone "100, 150, 180"
Initial Pruning speed up 2.0
Finetuning epoch 140
lr 0.01
weight decay 0.0005
milestone "80, 120"
Metanetwork num layer 8
hiddim 32
in node dim 5
in edge dim 9
node res ratio 0.01
edge res ratio 0.01
Meta Training lr 0.001
weight decay 0.0005
milestone "15"
pruner reg 10
Final Pruning metanetwork epoch 51
speed up 8.95
Finetuning After Metanetwork epoch 2000
lr 0.01
weight decay 0.0005
milestone "1850, 1950"
Finetuning After Pruning epoch 2000
lr 0.01
weight decay 0.0005
milestone "1850, 1950"
Appendix D ResNet50 on ImageNet
D.1 Equivalent Conversion between Network and Graph
We change ResNet50 into a graph with node featrues of 8 dimensions and edge features of 1, 9 or 49 dimensions. We employ 3 linear layers to project edge features of these 3 different dimensions into the same hidden dimension, enabling their integration into our graph neural network.
The node features consist of 4 features derived from the batch normalization parameters (weight, bias, running mean, and running variance) of the previous layer, along with 4 features from the batch normalization of the previous residual connection. If no such residual connection exists, we assign default values $[1,0,0,1]$ to the corresponding 4 features.
The edge features are constructed by simply flatten convolutional kernels into feature vectors (We have convolutional kernels of size $1\times 1$, $3\times 3$ or $7\times 7$). We treat all residual connections as convolutional layers with kernel size $1\times 1$. For example, consider two layers A and B connected via a residual connection, with neurons indexed as $1,2,3,\ldots$ in both layers. If the residual connection has no learnable parameters (i.e., it directly adds the input to the output), we represent the edge from $A_{i}$ to $B_{i}$ as a $1\times 1$ convolutional kernel with value 1. Different from section B, to reduce the VRAM footprint of the program, we build no edges between $A_{i}$ and $B_{j}$ ($i\neq j$). If the residual connection includes parameters (e.g., a downsample with a convolutional layer and batch normalization), we construct the edge features in the same way as for standard convolutional layers.
D.2 Meta Training
Data Models : We generate 3 models as our data models, among them, 2 are used for meta-training and 1 are used for validation (visualize the metanetwork). When generating each data model, we finetune them based on the pretrained weights for 30 epochs with learning rate 0.01 and milestone "10", then pruning with speed up 1.2920x, followed by a finetuning for 60 epochs with learning rate 0.01 and milestone "30". Finally, we can get a network with accuracy around 76.1%.
One Meta Training Epoch : We train with torch.nn.parallel.DistributedDataParallel (pytorch data parallel) across 8 gpus. Unlike in Section B and Section C, at the beginning of each epoch, we evenly distribute different data models across the GPUs. For example, since we have two data models, we load one on four GPUs and the other on the remaining four GPUs. At each iteration, we forward all eight models (replicated across the 8 GPUs) through the metanetwork to generate eight new models. We then feedforward a large batch of ImageNet data, evenly distributed through these eight new models, to compute the accuracy loss. We also use the parameters of the eight new networks to compute the sparsity loss. Then we backward the gradients from both two losses to update our metanetwork. Each training epoch consists of (ImageNet data num / big batch size) iterations, meaning that one full pass over the entire ImageNet dataset when computing the accuracy loss is considered as one epoch .
Training: We train our metanetwork with learning rate 0.01, milestone "4", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 13 as our final network for pruning.
D.3 Full Results (Including Error Bars)
All results are summarized in Table 10, where we repeat the pruning process 3 times. It is important to note that due to the design of our algorithm and implementation, we cannot prune the network to exactly the same speed up across different runs. For instance, when targeting a 2.3$\times$ speed up, the actual achieved speed up may be slightly larger, such as 2.31$\times$, 2.32$\times$, or 2.35$\times$.
The aggregated statistical results are presented in Table 11, where each value is reported in the form of mean(standard deviation).
Table 10: ResNet50 on ImageNet Full
Method Base Top-1(Top-5) Pruned Top-1($\Delta$) Pruned Top-5($\Delta$) Pruned FLOPs
DCP [77] 76.01%(92.93%) 74.95%(-1.06%) 92.32%(-0.61%) 55.6%
CCP [49] 76.15%(92.87%) 75.21%(-0.94%) 92.42%(-0.45%) 54.1%
FPGM [19] 76.15%(92.87%) 74.83%(-1.32%) 92.32%(-0.55%) 53.5%
ABCP [42] 76.01%(92.96%) 73.86%(-2.15%) 91.69%(-1.27%) 54.3%
DMC [11] 76.15%(92.87%) 75.35%(-0.80%) 92.49%(-0.38%) 55.0%
Random [38] 76.15%(92.87%) 75.13%(-1.02%) 92.52%(-0.35%) 51.0%
DepGraph [9] 76.15%(-) 75.83%(-0.32%) - 51.7%
ATO [62] 76.13%(92.86%) 76.59%(+0.46%) 93.24%(+0.38%) 55.2%
DTP [39] 76.13%(-) 75.55%(-0.58%) - 56.7%
ours 76.14%(93.11%) 76.17%(+0.03%) 92.94%(-0.17%) 56.9%
ours 76.14%(93.11%) 76.27%(+0.13%) 93.05%(-0.06%) 56.9%
ours 76.14%(93.11%) 76.15%(+0.01%) 92.95%(-0.16%) 56.9%
Table 11: ResNet50 on ImageNet Statistics
Pruned FLOPs (Speed Up) Pruned Top-1 Pruned Top-5
56.9%(2.32$\times$) 76.20%($\pm$0.05%) 92.98%($\pm$0.08%)
D.4 Computer Resources and Hyperparameters
See Table 12.
Table 12: ResNet50 on ImageNet Hyperparameters
Types Name Value
Compute Resources GPU NVIDIA A100 40G
parallel 8
Batch size small batch size 32 $\times$ 8
big batch size 32 $\times$ 8 $\times$ 200
Prepare Data Models data model num 3 ( 2 + 1 )
Generate Data Model epoch 30
lr 0.01
weight decay 0.0001
milestone "10"
Initial Pruning speed up 1.2920
Finetuning epoch 60
lr 0.01
weight decay 0.0001
milestone "30"
Metanetwork num layer 4
hiddim 11
in node dim 8
node res ratio 0.001
edge res ratio 0.001
Meta Training lr 0.01
weight decay 0.0005
milestone "4"
pruner reg 10
Final Pruning metanetwork epoch 13
speed up 2.3095
Finetuning After Metanetwork epoch 200
lr 0.01
weight decay 0.0001
milestone "120, 160, 185"
Finetuning After Pruning epoch 200
lr 0.01
weight decay 0.0001
milestone "120, 160, 185"
Appendix E More About Ablation Study
E.1 Visualize "Test Acc VS. Speed Up" Curves
As mentioned in section 4.5, we visualize the "Test Acc VS. Speed Up" curves of metanetworks in Resnet56 on CIFAR10 in figure 4, VGG19 on CIFAR100 in figure 5 and ResNet 50 on ImageNet in figure 6.
Figure 4: Resnet56 on CIFAR10 Visualization
Figure 5: VGG19 on CIFAR100 Visualization
Figure 6: ResNet50 on ImageNet Visualization
E.2 Finetuning after Metanetwork
(a) Finetune Different Epochs with Same Metanetwork
(b) Use Finetune to Get Better Performance
Figure 7: Finetune After Metanetwork (ResNet56 on CIFAR10)
As mentioned in section 3.5, our pruning pipeline can be summarized as:
$InitialPruning(Finetuning)\rightarrow Metanetwork(Finetuning)\rightarrow Pruning(Finetuning)$
In this pipeline, finetuning after pruning is a common way to improve the accuracy. However, finetuning after the metanetwork is first proposed by us and may raise questions for readers regarding its impact on performance—whether it is necessary or beneficial ? In this section, we aim to clarify the effects of finetuning after the metanetwork and demonstrate that it is both necessary and advantageous.
To illustrate this, we use the example of pruning ResNet56 on CIFAR10. In Figure 7(a), we visualize the "Test Accuracy vs. Speed Up" curves for models after the same metanetwork and after various amounts of finetuning. Without any finetuning after the metanetwork, the network’s accuracy is nearly zero (which is why there is no "finetune_0_epoch" line in the figure—it lies at the bottom). Directly proceeding to pruning and finetuning from such an unfinetuned network experimentally results in poor performance.
However, as we increase the number of finetuning epochs after the metanetwork, several interesting observations emerge:
  • Quickly Recover: The accuracy of our network after metanetwork quickly recovers after only very few finetuning, indicating that our metanetwork has the ability to preserve the accuracy (i.e. our accuracy loss in meta-training works)
  • Trend of Change: As the number of finetuning epochs increases, the flat portion of the curve becomes shorter, and the overall accuracy of this region improves.
When we feed the same network into metanetworks trained for different numbers of meta-training epochs and get several modified networks, then finetune them with different epochs to reach the same accuracy (as shown in Figure 7(b)), another important pattern emerges:
  • More Finetune Better Performance : A network obtained from a metanetwork with more meta-training epochs can typically achieve the same accuracy as one from a metanetwork with fewer meta-training epochs—but requires more finetuning. Importantly, the former usually results in better pruning performance. This is evident in Figure 7(b), where the green line (representing a higher meta-training epoch) exhibits a longer flat region, indicating improved robustness to pruning.
This suggests that the effectiveness of our method can be further improved by employing metanetworks trained for more meta-training epochs and applying additional finetuning afterward. In practice, however, we avoid excessive finetuning in order to keep the number of finetuning epochs within a reasonable range, allowing for multiple experimental trials.
Overall, all finetuning in our pipeline is necessary and effective.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The Abstract and Introduction sections accurately reflect the contribution and scope of the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of our work is discussed in the conclusion section (5).
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We propose no theoretical result that needs to be proved.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We open sourced our code and provide everything you need to reproduce our code in Appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provides a URL link to our released code. All information about datasets and how to run our code can be found together with our code.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In Ablation Study and Appendix we talk about all the details of our experiments and how to choose hyparameter.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We run each experiment 3 times and calculate the error bars, which you can see in Appendix.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report our computer resources for each experiment in Appendix
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: This paper conforms to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: There is no societal impact of the work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper has no risks and needs no safeguards.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The assets utilized in this paper are appropriately acknowledged. The associated license and terms of use have been clearly stated and duly adhered to.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We open sourced our code and model with clear explanations.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We involve no crowdsourcing subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We involve no crowdsourcing subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We didn’t involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
-Meta Pruning via Graph Metanetworks : A Meta Learning Framework for Network Pruning
+Meta Pruning via Graph Metanetworks : A Universal Meta Learning Framework for Network Pruning
 Abstract
-Network pruning, aimed at reducing network size while preserving accuracy, has attracted significant research interest.
-Numerous pruning techniques have been proposed over time.
-They are becoming increasingly effective, but more complex and harder to interpret as well.
-Given the inherent complexity of neural networks, we argue that manually designing pruning criteria has reached a bottleneck.
-To address this, we propose a novel approach in which we use a neural network to prune neural networks.
+We propose an entirely new meta-learning framework for network pruning.
+It is a general framework that can be theoretically applied to almost all types of networks with all kinds of pruning and has great generality and transferability.
+Experiments have shown that it can achieve outstanding results on many popular and representative pruning tasks (including both CNNs and Transformers).
+Unlike all prior works that either rely on fixed, hand-crafted criteria to prune in a coarse manner, or employ learning to prune ways that require special training during each pruning and lack generality.
+Our framework can learn complex pruning rules automatically via a neural network (metanetwork) and has great generality that can prune without any special training.
 More specifically, we introduce the newly developed idea of metanetwork from meta-learning into pruning.
@@
 In this paper, we first establish a bijective mapping between neural networks and graphs, and then employ a graph neural network as our metanetwork.
-We train a metanetwork that learns the pruning strategy automatically which can transform a network that is hard to prune into another network that is much easier to prune.
-Once the metanetwork is trained, our pruning needs nothing more than a feedforward through the metanetwork and the standard finetuning to prune at state-of-the-art.
-Our method achieved outstanding results on many popular and representative pruning tasks (including ResNet56 on CIFAR10, VGG19 on CIFAR100, ResNet50 on ImageNet).
+We train a metanetwork that learns the pruning strategy automatically and can transform a network that is hard to prune into another network that is much easier to prune.
+Once the metanetwork is trained, our pruning needs nothing more than a feedforward through the metanetwork and some standard finetuning to prune at state-of-the-art.
 Our code is available at https://github.com/Yewei-Liu/MetaPruning.
+Figure 1: For a pruning criterion, we use a metanetwork to change a hard to prune network into another easy to prune network for better pruning.
 1 Introduction
-Correspondence to Muhan Zhang
-With the rapid advancement of deep learning [33, 53], neural networks have become increasingly powerful.
+With the rapid advancement of deep learning (LeCun et al., 2015; Schmidhuber, 2015), neural networks have become increasingly powerful.
 However, this improved performance often comes with a significant increase in the number of parameters and computational cost (FLOPs).
 As a result, there is growing interest in methods to simplify these networks while preserving their performance.
-Pruning, which involves selectively removing certain parts of a neural network, has proven to be an effective approach [2, 17, 51].
+Pruning, which involves selectively removing certain parts of a neural network, has proven to be an effective approach (Cheng et al., 2024; He & Xiao, 2024; Reed, 1993).
 A key idea of a large number of previous pruning works is to remove the unimportant components of a neural network.
 To achieve this goal, various criteria that measure the importance of components of a neural network have been invented.
-Some criteria are based on norm or magnitude scores [15, 36, 55, 18].
-Some are based on how much the performance changes when certain parts of the network are removed or preserved [68, 46, 10, 66].
-Some others defined more complex criteria such as saliency and sensitivity [32, 34, 72].
-Once a pruning criterion is established, we can assess whether a model is easy or difficult to prune based on that criterion.
-For instance, using a simple $l_{2}$ norm criterion, a model whose parameters exhibit either very large or very small $l_{2}$ norms is clearly easier to prune compared to one with uniformly distributed $l_{2}$ norms.
-From this viewpoint, we can classify models as more or less suitable for pruning.
-This concept forms the foundation of our work.
-Previously, researchers have generally approached pruning improvement in two ways: either by refining the pruning criteria or by transforming the model from “bad” to “good” according to a given criterion.
+Some criteria are based on norm or magnitude scores (Han et al., 2016; Li et al., 2017b; Sun et al., 2024; He et al., 2018a).
+Some are based on how much the performance changes when certain parts of the network are removed or preserved (You et al., 2019b; Ma et al., 2023; Frankle & Carbin, 2019; Ye et al., 2020).
+Some others defined more complex criteria such as saliency and sensitivity (LeCun et al., 1989; Lee et al., 2019; Zhao et al., 2019).
+Once a pruning criterion is established, we can assess whether a model is easy or hard to prune based on that criterion.
+Technically, we consider a model easy to prune if we can prune a relatively large part of it with only little loss in accuracy.
+Otherwise, we consider it hard to prune.
+Prior researchers have generally approached pruning improvement in two ways: either by refining the pruning criteria or by transforming the model to be easier to prune according to a given criterion.
 The first approach is more straightforward and has been extensively studied in prior work.
@@
 However, several studies adopting this latter perspective have achieved promising results.
-Notably, sparsity regularization methods [61, 9, 60] exemplify this approach by encouraging models to become sparser, thereby facilitating more effective pruning.
-Our work follows the second approach and introduces a novel method inspired by sparsity regularization.
-We focus on this approach because, despite the many pruning criteria proposed, most lack solid theoretical foundations and have become overly complex.
-Therefore, in this work, we adopt a simple variant of the $l_{2}$ norm—the group norm from Fang et al.
-[9]—as our pruning criterion.
-Additionally, we bring the concept of metanetworks from meta-learning into pruning.
-A metanetwork is a network that takes another neural network as input and produces a modefied neural network as output.
-In our method, the metanetwork transforms a “bad” network into a “good” one, making it more amenable to pruning.
-Once trained, our metanetwork-based approach requires only a feedforward pass through the metanetwork and some standard finetuning to achieve state-of-the-art pruning results.
+Notably, sparsity regularization methods (Wen et al., 2016; Fang et al., 2023; Wang et al., 2021) exemplify this approach by encouraging models to become sparser, thereby facilitating more effective pruning.
+Our framework inherits the idea of the second way.
+A metanetwork is a network that takes another network as input and produces a modefied network as output.
+For any given pruning criterion, our framework aims at training a metanetwork that can transform a hard to prune network into another easier to prune network like figure 1.
 Our contributions can be summarized as follows:
-  • Introducing the idea of metanetwork in meta-learning into pruning for the first time.
-  • Proposing a novel pruning method that is simple and powerful.
-  • Achieving state-of-the-art performance on several popular pruning tasks.
-It is worth noting that this paper focuses solely on structural pruning, which does not rely on specialized AI accelerators or software to reduce memory usage and computational costs.
-This makes it more versatile and broadly applicable.
-However, all our methods can also be applied directly to unstructured pruning.
+  1. (1)
+    Introduce the idea of metanetwork from meta-learning into pruning for the first time.
+  2. (2)
+    Propose an entirely new universal pruning framework that is theoretically applicable to almost all types of networks with all kinds of pruning.
+  3. (3)
+    Present concrete implementations for our theoretical framework and achieve state-of-the-art performance on various practical pruning tasks.
+  4. (4)
+    Conduct further research on the flexibility and generality of our framework.
+Our method is entirely new and fundamentally different from all prior works, see Appendix A for a comparison with prior works to better understand our novel contributions and advantages.
 2 Related Work
@@
 Using neural networks to process other neural networks has emerged as an intriguing research direction.
-Early studies demonstrated that neural networks can extract useful information directly from the weights of other networks [58].
+Early studies demonstrated that neural networks can extract useful information directly from the weights of other networks (Unterthiner et al., 2020).
 We define a metanetwork as a neural network that takes another neural network as input and outputs either information about it or a modified network.
-Initial metanetworks were simple multilayer perceptrons (MLPs), while more recent designs incorporate stronger inductive biases by preserving symmetries [13] inherent in the input networks.
+Initial metanetworks were simple multilayer perceptrons (MLPs), while more recent designs incorporate stronger inductive biases by preserving symmetries (Godfrey et al., 2022) inherent in the input networks.
 Broadly, metanetwork architectures can be categorized into two perspectives.
-The weight space view applies specially designed MLPs directly on the model’s weights [71, 48, 73, 57, 74, 75, 56].
-In contrast, the graph view transforms the input network into a graph and applies graph neural networks (GNNs) to it [40, 30, 27].
-In this work, we adopt the graph view to build our metanetwork, leveraging the natural correspondence between graph nodes and neurons in a network layer, as well as the symmetries they exhibit [47].
+The weight space view applies specially designed MLPs directly on the model’s weights (Zaheer et al., 2017; Navon et al., 2023; Zhou et al., 2023a; Tran et al., 2024b; Zhou et al., 2023b; 2024; Tran et al., 2024a).
+And the graph view transforms the input network into a graph and applies graph neural networks (GNNs) to it (Lim et al., 2024; Kofinas et al., 2024; Kalogeropoulos et al., 2024).
+In this work, we adopt the graph view to build our metanetwork, leveraging the natural correspondence between graph nodes and neurons in a network layer, as well as the symmetries they exhibit (Maron et al., 2019).
 Given the maturity and effectiveness of GNNs, the graph metanetwork approach offers both elegance and practicality.
 2.2 Graph Neural Networks
-Graph neural networks (GNNs) are a class of neural networks specifically designed to operate on graph-structured data, where graphs consist of nodes and edges with associated features [63, 52, 29].
-GNNs have gained significant attention in recent years, with frameworks such as the message passing neural network (MPNN) [12] proving to be highly effective.
-In this paper, we employ the message passing framework, specifically using Principal Neighborhood Aggregation (PNA) [3] as the backbone architecture for our metanetwork.
-2.3 Sparsity Regularization-Based Pruning
+Graph neural networks (GNNs) are a class of neural networks specifically designed to operate on graph-structured data, where graphs consist of nodes and edges with associated features (Wu et al., 2021; Scarselli et al., 2009; Kipf & Welling, 2017).
+GNNs have gained significant attention in recent years, with frameworks such as the message passing neural network (MPNN) (Gilmer et al., 2017) proving to be highly effective.
+In this paper, we employ the message passing framework, specifically using Principal Neighborhood Aggregation (PNA) (Corso et al., 2020) as the backbone architecture for our metanetwork.
+2.3 Sparsity Regularization Based Pruning
 Networks that are inherently sparse tend to be easier to prune effectively.
 Prior research has explored various regularization techniques to encourage sparsity in neural networks to facilitate pruning.
-This idea of sparsity regularization has been widely adopted across many works [61, 14, 25, 44, 22, 43, 64, 65, 9, 60].
-Our method inherits this principle to some extent, as our metanetwork can be viewed as transforming the original network into a sparser version that is more amenable to pruning.
+This idea of sparsity regularization has been widely adopted across many works (Wen et al., 2016; Gordon et al., 2018; Huang & Wang, 2018; Lin et al., 2020c; He et al., 2017; Lin et al., 2019; Xia et al., 2022; 2024; Fang et al., 2023; Wang et al., 2021).
+Our method inherits this principle to some extent, as our metanetwork can be viewed as transforming the original network into a sparser version for better pruning.
 2.4 Learning to Prune & Meta Pruning
-Given the complexity of pruning, several previous works have leveraged neural networks [5, 22, 62, 1], reinforcement learning [50, 70, 23], and other techniques to automatically learn pruning strategies.
-Viewing pruning as a learning problem naturally leads to the idea of meta-learning, which learns “how to learn” [24].
-Some prior works have specifically applied meta-learning to pruning [45, 37].
-While our work also adopts a meta-learning framework, it differs fundamentally from these previous approaches by using a metanetwork.
-3 Method
-Our approach centers on training a metanetwork that learns the pruning strategy automatically and can transform a neural network that is difficult to prune into one that is easier to prune, thereby improving pruning effectiveness and efficiency.
-We begin by describing how our metanetwork transforms one network into another, including the conversion between network and graph, and the design of our metanetwork(a graph neural network).
-Next, we introduce the group $l_{2}$ norm criterion for meta-training and pruning.
-Then we detail the meta-training pipeline used to train the metanetwork.
-Finally, we present the complete pruning process.
-[…]
-3.1 Equivalent Conversion between Neural Network and Graph
-To enable a neural network to be processed by another network, we first need to represent the original network in a structured data format.
-In this work, we convert the neural network into a graph, drawing inspiration from prior works such as Kofinas et al.
-[30], Lim et al.
-[40], but with notable differences in implementation.
-To convert a network into a graph, we establish the correspondence between the components of a neural network and elements of a graph as follows:
-  • Node: Each neuron or convolutional channel in the network is represented as a node in the graph.
-  • Edge: An edge exists between two nodes if there is a direct connection between them in the original network.
-This includes connections between neurons or channels within linear or convolutional layers, as well as residual connections.
-  • Node Features: Node features comprise parameters associated with neurons, here we use the weight, bias, running mean, running var of batchnorm [26], including both batchnorms from previous adjacent layers and previous skip connection layers (if exist).
-  • Edge Features: Edge features encode the weights of connections between nodes.
-We treat linear connections and residual connections as special cases of convolutional connections with a kernel size of 1.
-For a $k\times k$ convolution between two channels, we linearize it into a $k^{2}$-dim edge feature vector.
-Since graph neural networks require all edge features to have the same dimension, we adopt one of two strategies to standardize them:
-  • Pad all convolutional kernels to the same size (e.g., the maximum kernel size in the network), then flatten them to form fixed-length edge feature vectors.
-  • Flatten convolutional kernels of varying sizes, then apply learned linear transformations to project them into a uniform feature dimension.
-By following these rules, we can generate a graph that corresponds to a given neural network.
-Conversely, the inverse transformation can be applied to convert a graph back into an equivalent network.
-In essence, we establish an equivalent conversion between networks and graphs.
-For a more intuitive understanding, see Figure 1.
-Figure 1: We visualize a two layer small CNN as example.
+Given the complexity of pruning, several previous works have leveraged neural networks (Dery et al., 2024; He et al., 2017; Wu et al., 2024; Chen et al., 2023), reinforcement learning (Rao et al., 2019; Yu et al., 2022; He et al., 2018b), and other techniques to automatically learn pruning strategies.
+Viewing pruning as a learning problem naturally leads to the idea of meta-learning, which learns “how to learn” (Hospedales et al., 2022).
+Some prior works have specifically applied meta-learning to pruning (Liu et al., 2019; Li et al., 2020).
+Our work also draws idea from meta-learning, but it learns in a way entirely different from all prior works.
+3 A Universal Meta-Learning Framework
+Our framework can be summarized in one single sentence (Figure 1):
+For a pruning criterion, we use a metanetwork to change a hard to prune network into another easy to prune network for better pruning.
+Pruning criterion: A pruning criterion measures the importance of specific components of a neural network.
+Typical examples include the $\ell_{1}$ norm and $\ell_{2}$ norm of weight vectors.
+During pruning, we compute an importance score for each prunable component according to the criterion, and remove components in ascending order of their scores.
+Easy or hard to prune: For a fixed pruning criterion, if we gradually prune the network and the accuracy drops quickly, we refer to it as hard to prune.
+Conversely, if the accuracy decreases slowly with progressive pruning, we call it easy to prune.
+Metanetwork: A metanetwork is a special type of neural network whose inputs and outputs are both neural networks.
+In contrast, standard neural networks typically take structured data such as images (CNNs), sentences (Transformers), or graphs (GNNs) as input.
+A metanetwork, however, takes an existing network as input and outputs another network, hence the term meta—indicating that it operates on neural networks themselves.
+4 A Specific Implementation Based on Graph Metanetworks
+The proposed framework is conceptually simple and straightforward. However, its practical implementation requires addressing several technical challenges:
+  1. (1)
+    How to build the metanetwork ?
+  2. (2)
+    How to train the metanetwork ?
+  3. (3)
+    How to design the pruning criterion ?
+We present our implementations step by step, dedicating one subsection to each challenge.
+The final subsection offers an overall summary of our method and a deeper understanding of it from a more holistic perspective.
+4.1 Metanetwork Design
+Drawing ideas from recent researches (section 2.1), we use a graph neural network (GNN) as our metanetwork.
+Please make sure you are familiar with the basic concepts of GNN before continuing.
+Our design does not cover any complex GNN theory, you only need to know what is a graph and the basic message passing algorithm.
+Figure 2: Conversion between networks and graphs: We visualize a two layer toy CNN as example.
 Our graph contains 6 nodes corresponding with 6 neurons in the network.
@@
 This means we treat it as with previous residual batchnorm of weight 1, bias 0, running mean 0, running var 1, performing the same as with no batchnorm.
-3.2 Graph metanetwork design
-After changing the input network into a graph, we use our metanetwork, a graph neural network, to process the input graph.
-We build our metanetwork based on the message passing framework [12] and PNA [3] architecture.
+Conversion between networks and graphs: To apply a metanetwork (GNN) on networks, we first need to establish a conversion between networks and graphs.
+Prior works such as Kofinas et al.
+(2024); Lim et al.
+(2024) have managed to change networks of almost all architectures (CNN, Transformer, RNN, etc.) into graphs.
+Here we show how we establish conversion between a ResNet (He et al., 2016) and a graph in our experiments as an example.
+This inherits the idea of Kofinas et al.
+(2024); Lim et al.
+(2024) but is quite different in implementation details.
+To convert a network into a graph, we establish the correspondence between the components of a neural network and elements of a graph as follows:
+(1) Node: Each neuron in fully connected layers or channel in convolutional layers in the network is represented as a node in the graph.
+(2) Edge: An edge exists between two nodes if there is a direct connection between them in the original network.
+This includes fully connected layers, convolutional layers, and residual connections.
+(3) Node Features: Node features comprise parameters associated with neurons, here we use the weight, bias, running mean, running var of batchnorm (Ioffe & Szegedy, 2015), including both batchnorms from previous adjacent layers and previous skip connection layers (if exist).
+(4) Edge Features: Edge features encode the weights of connections between nodes.
+We treat linear connections and residual connections as special cases of convolutional connections with a kernel size of 1.
+For a $k\times k$ convolution between two channels, we flatten it into a $k^{2}$-dim edge feature vector.
+Since graph neural networks here require all edge features to have the same dimension, we adopt one of two strategies to standardize them: (1) Pad all convolutional kernels to the same size (e.g., the maximum kernel size in the network) (2) Flatten convolutional kernels of varying sizes, then apply learned linear transformations to project them into the same size.
+During conversion, we simply ignored other components such as pooling layers.
+Although we have methods to explicitly convert all of them, our experiments show that omitting them already yields outstanding results.
+Therefore, it is acceptable to ignore them in practice.
+We also didn’t use any other techniques like positional embedding.
+By following these rules, we can generate a graph that corresponds to a given neural network.
+Conversely, the inverse transformation can be applied to convert a graph back into an equivalent network.
+In essence, we establish an equivalent conversion between networks and graphs.
+For a more intuitive understanding, see Figure 2.
+Metanetwork architecture: After establishing the conversion between networks and graphs, we can convert the input network into a graph, use metanetwork(GNN) to transform the graph, and finally convert the output graph back into a new network.
+Our metanetwork must be powerful enough to learn the transforming rules for better pruning.
+We build our metanetwork based on the message passing framework (Gilmer et al., 2017) and PNA (Corso et al., 2020) architecture.
+A brief overview of our architecture is provided below, the full description is available in Appendix B.1.
+We use $v_{i}$ to represent node $i$’s feature vector, and $e_{ij}$ to represent feature vector of the edge between $i$ and $j$.
+We consider undirected edges, i.e., $e_{ij}$ is the same as $e_{ji}$.
+Our metanetwork is as follows.
+First, all input node and edge features are encoded into the same hidden dimension.
+$v\leftarrow\text{MLP}_{NodeEnc}(v_{in}),$ (1)
+$e\leftarrow\text{MLP}_{EdgeEnc}(e_{in}).$ (2)
+Then they will pass through several message passing layers. In each layer, we update node and edge features using information from adjacent nodes and edges.
+$v_{i}$ $\leftarrow f_{v}(v_{i},v_{j}((i,j)\in E),e_{ij}((i,j)\in E))$ (3)
+$e_{ij}$ $\leftarrow f_{e}(v_{i},v_{j},e_{ij})$ (4)
+Here $f_{v}$ and $f_{e}$ are fixed calculating processes.
+Finally, we use a decoder to recover node and edge features to their original dimensions:
+$v_{pred}\leftarrow\text{MLP}_{NodeDec}(v),$ (5)
+$e_{pred}\leftarrow\text{MLP}_{EdgeDec}(e).$ (6)
+We multiply predictions by a residual coefficient and add them to the original inputs as final outputs:
+$v_{out}=\alpha\cdot v_{pred}+v_{in},$ (7)
+$e_{out}=\beta\cdot e_{pred}+e_{in},$ (8)
+where $\alpha$ and $\beta$ in practice are set to a small real numbers like $0.01$.
+This design enables the metanetwork to learn only the delta weights on top of the original weights, instead of a complete reweighting.
+4.2 Meta-Training
+Meta-training is the process we train our metanetwork (Figure 3).
+Figure 3: Meta-Training Pipeline: During each iteration, the origin network is converted into the origin graph, fed through the metanetwork to get a new graph and finally converted back into a new network.
+We calculate accuracy loss and sparsitly loss on the new network, then backpropagated the gradients to update the metanetwork.
+Data preparation: Meta-training data is comprised of two parts.
+The first is a set of neural networks to prune, which serve as the origin network in meta-training.
+We call them data models.
+The number of data models can be arbitrary small numbers like 1, 2, or 8, because our implementations aren’t sensitive to it (See more explanations in Section 4.4).The second is the traditional training datasets, CIFAR-10 for example, used to calculate the accuracy loss.
+One training iteration: We select one model from data models as the origin network.
+It is converted into its graph representation, passed through the metanetwork to get a new graph and converted back into a new network.
+We calculate accuracy loss and sparsity loss on this new network, and backpropagated the gradients to update the metanetwork.
+Two types of losses are used during meta-training.(1) Accuracy Loss: We feed the training data (e.g., Train set of CIFAR-10.
+Note that test set is not used here) into the new network and compute the cross-entropy loss based on the output predictions, ensuring the metanetwork doesn’t excessively disturbing the effective parts of the origin network.
+(2) Sparsity Loss: We calculate a regularization term on the new network to encourage it to be sparse, making it easier to prune.
+It is related to the design of pruning criterion (section 4.3)
+4.3 Pruning Criterion
+A valid pruning criterion consists of two components:(1) An importance score function: During pruning, we calculate the scores of all prunable components of a network, and prune them in ascending order of their scores.(2) A sparsity loss: It is a kind of loss that is calculated on the parameters of the network.
+Optimizing the network using this loss can make it easier to prune.
+We draw ideas from prior sparsity regularization based pruning methods(section 2.3), and design our pruning criterion based on them (especially Fang et al.
+(2023)).
+We will give a brief introduction below.
+For a rigorous mathematical definition and more details, please refer to appendix B.2.
+We only do structural pruning in all our experiments.
+A variant of the classical $\ell_{2}$ norm, the group norm is used as our importance score function.
+It inherits the idea of using $\ell_{2}$ norm to calculate a score, but calculate it on a pruning group level.
+Since structural pruning is performed on groups, traditional $l_{2}$ norm that treats nodes and edges in isolation is no longer viable.
+Group $l_{2}$ norm that takes a whole group into consideration has shown to be more effective.
+We calculate the sparisty loss by multiply our group norm scores with some coefficients.
+The choice of pruning criteria is highly flexible (see in our later experiments, section 6.1), which further demonstrates the generality of our framework.
+Although we conduct only structural pruning in all our experiments, all our methods can be directly applied to unstructured pruning as well.
+4.4 A Holistic Perspective
+Figure 4: Feed forward through metanetwork: when a network is fed forward through the metanetwork, every parameter in it gathers information from neighbour parameters and architectures.
+Intuitively, when a network is fed forward through the metanetwork, every parameter in it gathers information from neighbour parameters and architectures (Figure 4).
+The metanetwork automatically learns how to process the information from neighbours and change the weights of the network based on that for better pruning.
+And this has several important properties:
+Generality: Our metanetwork is a really small GNN compared to the origin network, which means each parameter can only see information from a limited local region.
+Due to the nature of GNN, parameters at any position in the network—whether at the beginning, middle, or end—must process information from neighbours in the same way.
+As a result, the metanetwork must learn a general strategy to adjust parameters based on neighbour information for better pruning.
+This explains why, in all our experiments, the pruning models and meta-training data models are completely different, but the performance is as good as the same.
+This also explains why our implementations are not sensitive to the number of data models.
+Even a single data model can be viewed as a collection of numerous data points sufficient to train the metanetwork.
+So there is no overfitting at all, and whether we are using 1, 2 or 8 data models during meta-training, the results remain the same.
+Natural transferability: Just as a GNN can be applied to graphs of varying sizes without modification, our metanetwork can be directly applied to networks of different scales.
+This property gives it natural transferability, as demonstrated in our experiments (section 6.2) where it successfully transfers across related datasets and network architectures.
+Universally applicability: Theoretically, our implementations can be applied to almost all types of networks with all kinds of pruning.
+Almost all types of networks can be converted into graphs (Kofinas et al., 2024; Lim et al., 2024), and the designs of the GNN and the pruning criteria are also broadly compatible.
+This makes our framework a universally applicable solution for network pruning.
+5 Experiments on Classical CNN Pruning Tasks
+5.1 Preliminaries
+$\underbrace{\text{Initial Pruning (Finetuning)}}_{\text{Optional}}\rightarrow\underbrace{\text{Metanetwork (Finetuning)}\rightarrow\text{Pruning (Finetuning)}}_{\text{Necessary}}$ (9)
+Definitions of the terms Speed Up and Acc vs. Speed Up Curve are in Appendix C.1. The pruning pipeline is described as equation 9, and full details are provided in Appendix C.2.
+Unlike prior learning to prune methods that require special training during each pruning, our pruning pipeline only requires a feed forward through metanetwork and standard finetuning, which is simple, general and effective.
+Once we have a metanetwork, we can prune as many networks as we want.
+All pruning models are completely different from meta-training data models.
+This demonstrates our metanetwork naturally has great transferability and no prior work has done something like this before.
+See Appendix A for more comparisons between our work and prior ones.
+5.2 Classical CNN Pruning Tasks
+We carry out our experiments on three most classical, popular, and representative image recognition tasks, including pruning ResNet56 (He et al., 2016) on CIFAR10 (Krizhevsky & Hinton, 2009), VGG19 (Simonyan & Zisserman, 2015) on CIFAR100 (Krizhevsky & Hinton, 2009) and ResNet50 (He et al., 2016) on ImageNet (Deng et al., 2009).
+See appendix D.1 for more general setups.
+Our method achieves outstanding results on all 3 tasks and is better than almost all prior works (Table 1).
+See full results compared with prior works in Table 6(ResNet56 on CIFAR10), Table 9(VGG19 on CIFAR100), Table 12(ResNet50 on ImageNet).
+See Appendix D.2 D.3 D.4 for implementation details on each tasks.
+Table 1: Results for 3 classical CNN pruning tasks including pruning (1) ResNet56 on CIFAR10, (2) VGG19 on CIFAR100 (3) ResNet50 on ImageNet.
+For full results compared with prior works, see (1) Table 6, (2) Table 9, (3) Table 12.
+Task Base Top-1(Top-5) Pruned Top-1($\Delta$) Pruned Top-5($\Delta$) Pruned FLOPs
+(1) 93.51% 93.64%(+0.13%) 65.6%
+(2) 73.65% 69.75%(-3.90%) 88.83%
+(3) 76.14%(93.11%) 76.13%(-0.01%) 92.78%(-0.33%) 57.2%
+5.3 Ablation Study
+How metanetwork works: We visualized the "Acc vs.
+Speed Up" curve of both origin network and network after feedforward through metanetwork and finetuning (Figure 5).
+They show that as the pruning speed up increases, the accuracy drops at a significantly slower rate after applying the metanetwork and finetuning.
+(a) ResNet56 on CIFAR10
+(b) VGG19 on CIFAR100
+(c) ResNet50 on ImageNet
+(d) ViT on ImageNet
+Figure 5: Metanetwork changes hard to prune network into easy to prune network.
+Trade-off between accuracy and speed up: During meta-training: as the number of training epochs increases, the metanetwork’s ability to make the network easier to prune becomes stronger.
+However, its ability to maintain the accuracy becomes weaker.
+As shown in Figure 6(a), with more training epochs, the "Acc VS.
+Speed Up" curve shifts downward, and the flat portion of the curve becomes longer.
+This characteristic allows us to adaptively meet different pruning requirements.
+If a higher level of pruning is desired and a moderate drop in accuracy is acceptable, a metanetwork trained for more epochs would be preferable.
+Conversely, if preserving accuracy is more important, a metanetwork trained for fewer epochs would be a better choice (Figure 6(b)).
+(a) Acc VS. Speed Up curves of metanetworks from different meta-training epochs
+(b) Feed network through metanetworks from different meta-training epochs, finetuning, then prune it progressively to find the maximum speed up that can maintain the accuracy above a certain threshold.
+Figure 6: Trends in meta-training. (ResNet56 on CIFAR10 as example)
+Finetuning after Metanetwork. See appendix E.1 for how number of finetuning epochs after metanetwork influence the pruning results.
+6 Experiments on Transferability and Flexibility
+6.1 Flexible Pruning Criterion
+Pruning criterion of our method is highly flexible.
+We build a series of reasonable pruning criterion (Appendix B.2).
+At first, we directly use MEAN REDUCE, MAX NORMALIZE, $\alpha=4$ as default, because this is the same as group norm in Fang et al.
+(2023).
+Later, we keep everything else the same, and try many different criteria, and find they all works well (Table 2).
+This demonstrates that our framework is robust and has many flexibilities.
+Table 2: Flexible pruning criterion: Pruning ResNet56 on CIFAR10, all experiments use the same origin network wtih Test Acc 93.51%, all criteria use MEAN REDUCE.
+We tried different values of Alpha and different ways of NORMALIZE, they all work well.
+Even in the Naive situation, where we use no NORMALIZE and no shrinkage strength (Alpha is 0), the results remain robust.
+NORMALIZE alpha Pruned Acc Pruned FLOPs Speed Up
+Default MAX 4 93.64% 65.64% 2.91
+Alpha MAX 0 93.37% 65.75% 2.92
+MAX 2 92.87% 68.15% 3.14
+MAX 6 93.36% 65.75% 2.92
+NORMALIZE MEAN 4 93.42% 65.52% 2.90
+NONE 4 93.08% 65.64% 2.91
+Naive NONE 0 93.04% 65.75% 2.92
+6.2 Transfer between Datasets and Architectures
+Section 4.4 mentioned that our metanetwork has natural transferability. Following experiments demonstrate that our metanetwork can transfer between similar datasets and network architectures.
+Transfer between datasets: 3 datasets, CIFAR10 (Krizhevsky & Hinton, 2009), CIFAR100 (Krizhevsky & Hinton, 2009) and SVHN (Goodfellow et al., 2014) are used.
+We train metanetwork and the to be pruned network on each of them, traverse through every possible combination, and use the metanetwork to prune the network.
+Results (Table 3) show that out metanetwork can transfer between similar datasets.
+See Appendix F.1 for full details
+Table 3: Transfer between datasets: All networks’ architecture is ResNet56.
+Columns represent the training datasets for the metanetwork, and rows represent the training datasets for the to be pruned network.
+“None” indicates using no metanetwork.
+Results with metanetwork is obviously better than no metanetwork (The only exception is when training datasets for the to be pruned network is SVHN, and we guess this is because the dataset SVHN itself is too easy).
+[…]
+CIFAR10 93.35 92.47 92.87 91.28
+CIFAR100 69.97 70.16 69.25 68.91
+SVHN 96.79 96.50 96.86 96.78
+Transfer between architectures: 2 architectures, ResNet56 and ResNet110 (He et al., 2016) are used.
+We train metanetwork and the to be pruned network on each of them, traverse through every possible combination, and use the metanetwork to prune the network.
+Results (Table 4) show that our metanetwork can transfer between similar datasets.
+See Appendix F.2 for full details
+Table 4: Transfer between architectures.
+All training dataset is CIFAR10.
+Columns represent the architectures used for training the metanetwork, and rows represent the architecures of the to be pruned network.
+“None” indicates using no metanetwork.
+All results with metanetwork is obviously better than no metanetwork.
+[…]
+ResNet56 93.40 92.81 92.08
+ResNet110 93.04 93.38 92.40
+Possible Future Use: For large-scale networks, the metanetwork can be pretrained on smaller architectures of similar design.
+And when the original training dataset is unavailable or excessively large, a related dataset may be used for pretraining.
+This transferability substantially enhances the practicality of the framework.
+7 Experiments on Transformers
+As mentioned in section 4.4, our framework is theoretically applicable to almost all types of networks with all kinds of pruning.
+Here we expand our implementations to another widely used arthitecture–transformer (Vaswani et al., 2017).
+More specifically, we pruned the vision transformer (ViT) from Dosovitskiy et al.
+(2021) that is trained on ImageNet.
+See Appendix G for full experiments, including how we convert transformer into graph and conduct further meta-training and pruning.
+8 Conclusion
+We propose an entirely new meta-learning framework for network pruning.
+For a pruning criterion, we use a metanetwork to change a hard to prune network into another easy to prune network for better pruning.
+This is a general framework that can be theoretically applied to almost all types of networks with all kinds of pruning.
+We present practical implementations of our framework and achieve outstanding results.
+Further analysis and experiments show that our framework has natural generality, flexibility, and transferability.
+9 Reproducibility statement
+Our code is available at https://anonymous.4open.science/r/MetaPruning together with all the guides to reproduce our experiments.
+Appendix A Comparison with prior works
+Our framework is entirely new and fundamentally different from all prior works.
+In this section, we will compare our method with earlier ones to help you better understand the relationship between our method and prior ones, and realize our unique innovations and advantages.
+A.1 General Comparison
+Previous pruning approaches can be broadly categorized into fixed pruning methods and learning to prune methods.
+Our approach falls into the learning to prune category but learns in a completely different way.
+For clarity, we provide a detailed comparison in Table 5 to help you better understand how our work relates to previous studies.
+Table 5: Our framework is entirely new and has many advantages over previous works
+Fixed pruning Learning to prune Ours
+Definition Pruning with fixed hand-crafted algorithms. Using neural network learning techniques to learn how to prune. Also learning to prune, but learning in a completely different way.
+Performance Bad. Networks are complex and hand-crafted algorithms are quite limited Good. Good (almost best according to our experiments).
+Cost Low.
+No special extra training needed.
+High.
+Need special extra training during each pruning.
+Need special extra training before pruning.
+But once the training is done, can prune as many networks as we want without any special extra training.
+Potential Has already been well-studied and commonly used.
+The pruning process is too tricky and costly.
+Can only be used in a very specific situation and lack of generality.
+The pruning process is relatively more efficient.
+Can theoretically be used in almost all situations and has great generality.
+A.2 Computational and Memory Costs
+One important question for almost all previous meta-learning approaches is the computational and memory costs are too large.
+We provide some estimate of memory and computational costs for scaling our methods to larger models and compare with two classic pruning methods–pruning at initalization methods (cheap pruning, hard to training and low accuracy like Lee et al.
+(2019); Wang et al.
+(2022)) and iterative magnitude pruning (costly, pruning during training and high performance like Frankle & Carbin (2019); Molchanov et al.
+(2019)).
+The results show that our method achieves outstanding results while uses relatively low costs.
+For convenience in expression, we refer to pruning at initialization as “init pruning”, iterative magnitude pruning as “iter pruning” and our method as “meta pruning”.
+For all usual networks which satisfy Edge Number >> Neuron Number , the memory and computational cost are all $O(E)$ ($E=$ Edge Number).
+This is obvious for init and iter pruning.
+For meta pruning, our metanetwork doesn’t scale with the network.
+So all computation and memory cost is it still $O(E)$.
+PS: The reason why our metanetwork doesn’t scale with the network is that it only learns how to change weights based on local architectures and weights (the range of “local” is fixed for network of different sizes).
+In all our experiments, it is a small network compared to the to be pruned network, and it works well.
+So we have the confidence that it shouldn’t and doesn’t need to scale with the networks.
+While all methods share the same asymptotic cost, their constant factors differ significantly.
+Computational cost breaks down as:
+$\text{Compu Cost}=\text{Training cost}+\text{Pruning cost}+\text{Other cost}$
+  • Init pruning: Low training and pruning costs, negligible other cost.
+  • Iter pruning: High training cost (initial training + multiple finetuning steps) and high pruning cost (multiple pruning rounds), no extra cost.
+  • Meta pruning: Moderate training cost (initial training + 3 finetuning steps), moderate pruning cost (2 pruning steps), plus a small other cost consisting of metanetwork feedforward and amortized meta-training cost per network.
+We estimate the relative magnitude in the following table:
+Computation Training cost Pruning cost Other cost
+Init 1T 1P 0
+Iter >10>10T >5>5P 0
+[…]
+Where:
+T: Unit training cost
+P: Unit pruning cost
+[…]
+N: Number of networks pruned; meta-training cost is amortized over N (Once we get a metanetwork by meta-training, we can use it to prune as many networks as we want)
+A: Meta-training cost. A is estimated 5–20.
+Memory cost depends on both the amount of memory used and the duration of usage:
+Method Amount Duration All
+Init Small Short Low
+Iter Medium Long Large
+Meta Large for a very short time, Medium Medium but requires
+Small rest of the time high memory capacity
+  • Init pruning: Low memory usage for a short time.
+  • Iter pruning: Medium memory usage for a very long time.
+  • Meta pruning: The “feed forward through metanetwork” process requires large memory usage but takes a very short time.
+For the rest of the time it uses little memory.
+It is faster than iter pruning but slower than init pruning.
+While the average memory cost is moderate, peak usage demands high memory capacity.
+Meta-training requires large memory usage for a long time.
+But its cost can be amortized into each pruning like we mentioned before.
+Take pruning resnet50 or ViT on ImageNet as example (which is already quite large model and dataset), NVIDIA A100 with 80GiB VRAM is enough for meta-training and feedforward through metanetwork and the rest training and finetuning can be done on NVIDIA RTX 4090 with 24 GiB VRAM.
+When we don’t have enough memory capacity, we can also change the batch size and use more time to make up for the lack of our memory capacity.
+In summary, our method achieves outstanding results while uses relatively low costs.
+A.3 Generality
+Almost all previous meta-learning based pruning approaches, such as Liu et al.
+(2019); Li et al.
+(2020); Wu et al.
+(2024), are tailored to a specific network and thus lack generality.
+Because of this they require specialized training for each pruning instance.
+Unlike prior methods, our metanetwork learns a universal rule for adapting weights based on local information, thereby enabling more effective pruning without relying on layer-specific or architecture-specific heuristics.
+We need no special training during each pruning.
+Once our metanetwork is trained, it can be used to prune as many networks as we want, and even transfer between datasets and architectures.
+Appendix B Framework implementation details
+B.1 Metanetwork(GNN) Architecture
+We build our metanetwork based on the message passing framework (Gilmer et al., 2017) and PNA (Corso et al., 2020) architecture.
+Notation:
+$a\leftarrow b$ : $\quad\text{: assignment/update, overwrite }a\text{ with the value of }b$ (10)
+$X\odot Y$ : $\quad\text{: Hadamard/elementwise product, }(X\odot Y){i}=X_{i}Y_{i}$ (11)
 We use $v_{i}$ to represent node $i$’s feature vector, and $e_{ij}$ to represent feature vector of the edge between $i$ and $j$.
@@
 $e_{pred}\leftarrow\text{MLP}_{EdgeDec}(e).$ (21)
-We multiply our predictions by a residual coefficient and add them to the original inputs as our outputs:
+We multiply predictions by a residual coefficient and add them to the original inputs as our final outputs:
 $v_{out}=\alpha\cdot v_{pred}+v_{in},$ (22)
@@
 This design enables the metanetwork to learn only the delta weights on top of the original weights, instead of a complete reweighting.
-3.3 Group $l_{2}$ Norm Criterion
-We use group $l_{2}$ norm from Fang et al.
-[9] as our pruning criterion.
-The design philosophy of group $l_{2}$ norm is to measure the sparsity of pruning groups.
-In structural pruning, we only prune a group of elements in a network.
-For example, if we prune a neuron, then all edges connected to the neuron should also be pruned, which we say that they are in a group.
-A more detailed explanation of group can be found in Fang et al.
-[9].
-Since structural pruning is performed on groups, traditional $l_{2}$ norm that treats nodes and edges in isolation is no longer viable.
-Group $l_{2}$ norm that takes a whole group into consideration has shown to be more effective.
-Given a parameter group $g=\{w_{1},w_{2},\ldots,w_{|g|}\}$. For each parameters $w$ with $K$ prunable dimensions indexed by $w[k]$, we measure the importance of the k-th prunable dimension by :
-$I_{g,k}=\sum_{w\in g}\|w[k]\|^{2}$ (13)
+B.2 Pruning Criterion
+B.2.1 DepGraph and Torch-Pruning
+DepGraph (Fang et al., 2023) proposes a general sparsity regularization based structural pruning framework that can be applied to a wide range of neural network architectures, including CNNs, RNNs, GNNs, Transformers, etc.
+Alongside the paper, the authors released Torch-Pruning, a powerful Python library that enables efficient structural pruning for most modern architectures.
+In our work, the pruning criterion is designed based on the methodology of DepGraph (Fang et al., 2023), and all pruning operations are implemented using Torch-Pruning.
+Figure 7: A picture from (Fang et al., 2023). It shows how we group parameters together in structural pruning.
+B.2.2 Structural pruning
+In all our experiments, we only do structural pruning.
+Although we did not conduct experiments with unstructured pruning, our implementations are directly applicable to that setting as well.
+Structural pruning involves removing parameters in predefined groups, so that the resulting pruned model can be used as a standalone network—without relying on the original model with masks, nor requiring specialized AI accelerators or software to realize reductions in memory footprint and computational cost.
+In structural pruning, a pruning group consists of all parameters that must be pruned together to maintain network consistency.
+For example, in Figure 7(a), if an input channel of convolution layer $f_{4}$ is pruned, the corresponding channel in batch normalization (BN) layer $f_{2}$ and the output channel of convolution layer $f_{1}$ must be pruned as well.
+Furthermore, due to the presence of a residual connection, the corresponding channels in BN layer $f_{5}$ and the output channel of convolution layer $f_{4}$ must also be removed.
+All of these channels together form a single pruning group.
+We define the number of prunable dimensions of a group as the size of the input channel dimension of Conv $f_{4}$ (which is equivalent to the corresponding dimensions in BN $f_{2}$, the output channel of Conv $f_{1}$, etc.)
+For a more comprehensive theory to find all pruning groups in structural pruning, refer to the DepGraph (Fang et al., 2023) paper (section 3.1 & 3.2).
+B.2.3 A importance score function
+Given a parameter group $g=\{w_{1},w_{2},\ldots,w_{|g|}\}$ with K prunable dimensions indexed by $w_{t}[k]$ ($t\in\{1,2...|g|\}$), the score of the k th prunable dimension in group g is written is $I_{g,k}$, we introduce a general way to generate a series types of importance scores.
+The way has two key concepts, REDUCE and NORMALIZE.
+We first use REDUCE to reduce scores of parameters in the same group and in the same prunable dimension into one score.
+$I_{t,k}=\mathop{\text{REDUCE}}\limits_{t:t\in\{1,2...|g|\}}(|w_{t}[k]|^{p})$ (24)
+Here p is a hyperparameter and we usually set it to 2, REDUCE can be (but is not limited to) the following:
+  1. (1)
+    MEAN: $(w_{1}+w_{2}+...+w_{|g|})/|g|$
+  2. (2)
+    FIRST: $w_{1}$
+Then we use NORMALIZE to normalize scores in the different prunable dimensions of the same pruning group.
+$\hat{I}_{g,k}=I_{g,k}/\mathop{\text{NORMALIZE}}\limits_{k:k\in\{1,2...K\}}(I_{g,k})$ (25)
+Here NORMALIZE can be (but is not limited to) the following:
+  1. (1)
+    NONE: 1 (use no normalize).
+  2. (2)
+    MEAN: $(I_{g,1}+I_{g,2}+...+I_{g,K})/K$
+  3. (3)
+    MAX: The maximum amoung $I_{g,1},I_{g,2}...I_{g,K}$
+We calculate scores of all prunable dimensions in all groups in a network, rank them and prune according to their scores from lower to higher.
+B.2.4 A sparsity loss
 During training, our sparsity loss is defined as :
-$\mathcal{R}(g,k)=\sum_{k=1}^{K}\gamma_{k}\cdot I_{g,k}=\sum_{k=1}^{K}\sum_{w\in g}\gamma_{k}\|w[k]\|^{2}$ (14)
+$\mathcal{R}(g,k)=\sum_{k=1}^{K}\gamma_{k}\cdot\hat{I}_{g,k}$ (26)
 Where $\gamma_{k}$ refers to a shrinkage strength applied to the parameters to modify the gradients for better training. Defined as :
-$\gamma_{k}=2^{\alpha(I_{g}^{\max}-I_{g,k})(I_{g}^{\max}-I_{g}^{\min})}$ (15)
-And we use $\alpha=4$ in all our experiments.
-When pruning, we calculate the normalized relative score of all prunable dimensions in all groups
-$\hat{I}_{g,k}=N\cdot I_{g,k}/\sum\{\text{TopN}(I_{g})\}$ (16)
-where $N$ is a hyperparameter that chooses the top $N$ importance scores in the group $g$ to normalize every prunable dimension $k$’s score $I_{g,k}$.
-We prune unimportant dimensions of groups globally from lower to higher according to the normalized relative score.
-Figure 2: Meta-Training Pipeline.
-3.4 Meta-Training Pipeline
-The goal of training our metanetwork to enable it to transform a hard-to-prune network into an easy-to-prune one, referred to as meta-training.
-This section details our meta-training pipeline, which is illustrated in Figure 2.
-To prepare for meta-training, first we need the training data.
-Meta-training data is comprised of two parts.
-The first is a set of neural networks to prune, typically between 2 to 8, which serve as the origin network in meta-training.
-The second is the traditional training datasets, such as CIFAR-10, CIFAR-100, or ImageNet, used to calculate the accuracy loss.
-During each training iteration, we select one model from data models as the origin network.
-The origin network is then converted into its graph representation.
-This graph is passed through the metanetwork, which outputs a new graph.
-We subsequently convert this new graph back into a new network.
-Two types of losses are computed on the new network:
-1) Accuracy Loss: We feed the training data into the new network and compute the cross-entropy loss based on the output predictions, ensuring the new network to have high accuracy after finetuning.
-2) Sparsity Loss: We apply the group $l_{2}$ norm (as defined in 14) on the new network to encourage structured sparsity, making it more amenable to pruning.
-The gradients from the loss are backpropagated through the metanetwork to update its parameters.
-The accuracy loss ensures that the metanetwork preserves the performance of the original network, while the sparsity loss encourages the metanetwork to generate a network with group sparsity that facilitates pruning.
-After meta-training, our metanetwork can generate a transformed network that has almost no loss in accuracy after finetuning, and is significantly easier to prune.
-Empirically, we observe that the metanetwork performs much better on hard-to-prune origin networks.
-However, most networks are initially easy to prune—for example, ResNet-56 on CIFAR-10 can be pruned with a 1.3× speedup and finetuned without accuracy loss.
-Therefore, in all our experiments (both during meta-training and pruning), we first prune the network slightly—without any accuracy degradation—to make it harder to prune.
-This pre-pruning step allows us to fully leverage the capability of the metanetwork.
-3.5 Pruning a Target Network
-Our pruning pipeline only involves a feedforward through metanetwork and finetuning. It can be summarized as:
-$InitialPruning(Finetuning)\rightarrow Metanetwork(Finetuning)\rightarrow Pruning(Finetuning)$
-As discussed in 3.4, we first do initial pruning to make our network harder to prune while maintaining the accuracy.
-Then we feed the hard-to-prune network through the metanetwork to transform it into a much easier-to-prune one, which also has almost no accuracy loss after finetuning.
-In the end, we perform the final pruning to reduce its size, and the final finetuning to evaluate the performance.
-4 Experiments
-4.1 Setup
-We carry out our experiments on three most classical, popular, and representative image recognition tasks, including resent56 [16] on CIFAR10 [31], VGG19 [54] on CIFAR100 [31] and resnet50 [16] on ImageNet [4].
-See appendix A for more general setups.
-[…]
-4.2 ResNet56 on CIFAR10
-Table 1: ResNet56 on CIFAR10 (see full table in appendix 4)
-Method Base Pruned $\Delta$ Acc Pruned FLOPs Speed Up
-FPGM [19] 93.59 92.93 -0.66 52.6% 2.11
-FPC [21] 93.59 93.24 -0.35 52.9% 2.12
-DMC [11] 93.62 92.69 -0.93 50.0% 2.00
-GNN-RL [70] 93.49 93.59 0.10 54.0% 2.17
-DepGraph w/o SL [9] 93.53 93.46 -0.07 52.6% 2.11
-DepGraph with SL [9] 93.53 93.77 0.24 52.6% 2.11
-ATO [62] 93.50 93.74 0.24 55.0% 2.22
-Meta-Pruning (ours) 93.51 94.08 0.57 56.5% 2.30
-C-SGD [7] 93.39 93.44 0.05 60.8% 2.55
-Greg-1 [60] 93.36 93.18 -0.18 60.8% 2.55
-Greg-2 [60] 93.36 93.36 0.00 60.8% 2.55
-DepGraph w/o SL [9] 93.53 93.36 -0.17 60.2% 2.51
-DepGraph with SL [9] 93.53 93.64 0.11 61.1% 2.57
-ATO [62] 93.50 93.48 -0.02 65.3% 2.88
-Meta-pruning (ours) 93.51 93.69 0.18 65.6% 2.91
-Meta-pruning (ours) 93.51 93.25 -0.26 66.9% 3.02
-Table 1 shows our results of pruning ResNet56 [16] on CIFAR10 [31].
-Our algorithm outperforms all previous works.
-We can prune 65.6% FLOPs (i.e.
-2.91x Speed Up) without any loss of accuracy, which is quite amazing.
-See Appendix B for more results and implementation details.
-4.3 VGG19 on CIFAR100
-Table 2: VGG19 on CIFAR100 (see full table in appendix 7)
-Method Base Pruned $\Delta$ Acc Pruned FLOPs Speed Up
-OBD [59] 73.34 60.66 -12.68 83.58% 6.09
-EigenD [59] 73.34 65.18 -8.16 88.64% 8.80
-Greg-2 [60] 74.02 67.75 -6.27 88.69% 8.84
-DepGraph w/o SL [9] 73.50 67.60 -5.44 88.73% 8.87
-DepGraph with SL [9] 73.50 70.39 -3.11 88.79% 8.92
-Meta-Pruning (ours) 73.65 70.06 -3.59 88.85% 8.97
-Table 2 shows our results of pruning VGG19 [54] on CIFAR100 [31].
-Our algorithm reaches 70.06% pruned accuracy with 88.85% FLOPs pruned (i.e.
-8.97x Speed Up).
-While our pruning speed up is larger than all previous works, our pruned accuracy is only a little bit lower than DepGraph with SL and higher than all other works.
-See Appendix C for more results and implementation details.
-4.4 ResNet50 on ImageNet
-Table 3: ResNet50 on ImageNet (see full table in appendix 10)
-Method Base Top-1(Top-5) Pruned Top-1($\Delta$) Pruned Top-5($\Delta$) Pruned FLOPs
-DCP [77] 76.01%(92.93%) 74.95%(-1.06%) 92.32%(-0.61%) 55.6%
-CCP [49] 76.15%(92.87%) 75.21%(-0.94%) 92.42%(-0.45%) 54.1%
-FPGM [19] 76.15%(92.87%) 74.83%(-1.32%) 92.32%(-0.55%) 53.5%
-ABCP [42] 76.01%(92.96%) 73.86%(-2.15%) 91.69%(-1.27%) 54.3%
-DMC [11] 76.15%(92.87%) 75.35%(-0.80%) 92.49%(-0.38%) 55.0%
-Random [38] 76.15%(92.87%) 75.13%(-1.02%) 92.52%(-0.35%) 51.0%
-DepGraph [9] 76.15%(-) 75.83%(-0.32%) - 51.7%
-ATO [62] 76.13%(92.86%) 76.59%(+0.46%) 93.24%(+0.38%) 55.2%
-DTP [39] 76.13%(-) 75.55%(-0.58%) - 56.7%
-ours 76.14%(93.11%) 76.17%(+0.03%) 92.94%(-0.17%) 56.9%
-Table 3 shows our results of pruning ResNet50 [16] on ImageNet[4].
-Our algorithm reaches 76.17% top-1 and 92,95% top-2 accuracy with 56.9% FLOPs pruned.
-With the largest pruned FLOPs, our pruned accuracy is higher than almost all previous works (including DTP [39]).
-The only exception is that our accuracy is a bit lower than ATO [62], but our pruned FLOPs is much larger (1.7%) than it.
-See appendix D for more results and implementation details.
-4.5 Ablation Study
-How metanetwork works.
-To help readers visually grasp the effect of metanetwork, we visualized the "Test Acc VS.
-Speed Up" curve of both origin network and network after feedforward through metanetwork and finetuning .
-We generate the curve by prune the network little by little.
+$\gamma_{k}=2^{\alpha\frac{\sqrt{I_{g}^{\max}}-\sqrt{I_{g,k}}}{\sqrt{I_{g}^{\max}}-\sqrt{I_{g}^{\min}}}}$ (27)
+Here $\alpha$ can be but is not limited to:
+  1. (1)
+    4: What we use as default in all our experiments.
+  2. (2)
+    0: Same as using no shrinkage strength.
+In sparsity loss, we treat $\gamma_{k}$ and NORMALIZE in the denominator of $\hat{I}_{g,k}$ simply as constants, which means gradients aren’t backpropagated through them.
+Gradients only back propagated from $\mathcal{R}(g,k)$ to $I_{g,k}$ to the parameters of the network.
+In our code, rather than calculate the sparsity loss, we directly modify the gradients of the parameters of the network, which get the same results.
+B.2.5 Kinds of pruning criterion
+By choosing different ways of REDUCE and NORMALIZE (appendix B.2.3) and different $\alpha$ (appendix B.2.4), we can make different pruning cirteria in a uniform framework.
+In most of our experiments, we use MEAN REDUCE, MAX NORMALIZE, $\alpha=4$, as our default pruning criterion, and we name it group $\ell_{2}$ norm max normalizer.
+We also tried different criteria in our experiments and found they are almost all effective, which further shows the flexibility and effectiveness of our framework.
+B.2.6 Pay Attention
+Though we draw ideas from Fang et al.
+(2023) and use their Torch-Pruning libirary, we find their paper is inconsistent with their code.
+We wrote appendix B.2.3 & B.2.4 strictly based on their code, which is the Torch-Pruning library.
+If you read the Fang et al.
+(2023) paper section 3.3, you may find our description a little different from theirs, because their paper is inconsistent with their code, and we are in line with their code.
+Appendix C Experimental Preliminaries
+C.1 Terminologies
+Speed Up: In line with prior works, we define $\textbf{Speed Up}\triangleq\textbf{Origin FLOPs }/\textbf{ Pruned FLOPs}$.
+It reflects the extent to which our pruning reduces computation, thereby accelerating the network’s operation.
+Acc vs.
+Speed Up Curve: We generate the curve by prune the network little by little.
 After each pruning step, we evaluated the model’s performance on the test set (no finetuning in this step) and recorded a data point representing the current speed up and corresponding test accuracy.
 Connecting these points forms a curve that illustrates how accuracy changes as the speed up increases.
-The "Test Acc VS.
-Speed Up" curve of ResNet on CIFAR10 can be seen in figure 3(a).
-Similar curves of VGG19 on CIFAR100 and ResNet50 on ImageNet can be seen in appendix E.1.
-As shown in the "Test Accuracy vs.
-Speed Up" curves, as the pruning speed up increases, the accuracy drops at a significantly slower rate after applying the metanetwork and finetuning.
-Take figure 3(a) as an example (this is exactly the metanetwork we use in 4.2).
-For the original network, the accuracy decreases rapidly and nearly collapses before reaching a speed up of 1.50x, suggesting that directly using group l2 norm criterion 3.3 to prune the origin network removes many important parts of the network (i.e.
-the origin network is hard to prune under this criterion).
-However, after metanetwork and finetuning, pruning to 2.0x speed up without finetuning causes almost no drop in accuracy, suggesting that using group l2 norm criterion on this network can efficiently distinguish the unimportant parts of this network (i.e.
-the network after metanetwork and finetuning is much easier to prune under this criterion).
-This supports our claim in 3 that metanetwork can transform a neural network that is difficult to prune into one that is much easier to prune.
-(a) Origin VS After Metanetwork
-(b) Impact Of Training Epochs On Metanetworks
-Figure 3: Visualization of how metanetworks effect the networks (ResNet56 on CIFAR10)
-Trade-off between accuracy and speed up.
-We observe an interesting phenomenon during meta-training: as the number of training epochs increases, the metanetwork’s ability to make the target network easier to prune becomes stronger.
-However, this comes at the cost of its ability to maintain the accuracy.
-As shown in figure 3(b), with more training epochs, the "Test Accuracy vs.
-Speed Up" curve shifts downward, and the flat portion of the curve becomes longer.
-This characteristic allows us to adaptively meet different pruning requirements.
-If a higher level of pruning is desired and a moderate drop in accuracy is acceptable, a metanetwork trained for more epochs would be preferable.
-Conversely, if preserving accuracy is more important, a metanetwork trained for fewer epochs would be a better choice.
-Finetuning after Metanetwork. See appendix E.2 for our study on the impact of the number of finetuning epochs after metanetwork on performance.
-5 Conclusion
-In this paper we proposes a brand new meta-learning framework for network pruning.
-For the first time we introduce the concept of metanetwork (a network that takes another network as input and produces a modified network as output) into pruning.
-The metanetwork is trained to automatically learn effective pruning strategies.
-After training, it can take a network that is hard to prune as input and generate a transformed network that is much easier to prune.
-During pruning, our method needs nothing more than a feedforward through metanetwork and trivial finetuning to prune at state-of-the-art.
-Advantages Over Previous Works.
-Our metanetwork is capable of considering the weights of multiple nearby layers, which sets it apart from most previous approaches.
-Unlike earlier methods that either focus only on local weights or incorporate more layers but rely on simple, hand-crafted relationships between them, our metanetwork automatically learns complex relationships among neighboring layers.
-These relationships are too intricate to be expressed by simple human-defined rules.
-So we use metanetwork to learn it automatically during training.
-This can explain why our work outperforms almost all previous works,
-Limitations And Future Work.
-Our work also has several limitations.
-First, converting networks into graph representations requires GPUs with high VRAM, making it challenging to apply our method to large and densely parameterized models.
-Second, our current approach can’t be directly applied to networks with different architectures, such as Transformers or RNNs without appropriate modifications.
-However, these limitations also present opportunities for future research.
-We think the following directions are promising and worth exploring: (1) Extending our method to accommodate diverse architectures, including Transformers, diffusion models, and others.
-(2) Investigating alternative approaches to constructing the metanetworks.
-(3) Applying similar meta-learning frameworks to a broader range of tasks.
-Acknowledgments and Disclosure of Funding
-Use unnumbered first level headings for the acknowledgments.
-All acknowledgments go at the end of the paper before the list of references.
-Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work).
-More information about this disclosure can be found at: https://neurips.cc/Conferences/2025/PaperInformation/FundingDisclosure.
-Do not include this section in the anonymized submission, only in the final paper.
-You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
-Appendix A General Experiment Setup
-A.1 General Settings
+C.2 Pruning Pipeline
+Our pruning pipeline can be summarized as:
+$\underbrace{\text{Initial Pruning (Finetuning)}}_{\text{Optional}}\rightarrow\underbrace{\text{Metanetwork (Finetuning)}\rightarrow\text{Pruning (Finetuning)}}_{\text{Necessary}}$ (28)
+Initial Pruning: Empirically, we observe that the metanetwork performs better on hard to prune origin networks.
+However, many networks are initially easy to prune.
+For instance, ResNet56 on CIFAR10 can be pruned with a $1.3\times$ speedup and then finetuned without any accuracy loss.
+Therefore, unless stated otherwise, we apply the initial pruning step in all our experiments—both when generating meta-training data models and during pruning.
+In this step, we slightly prune the original network with a fixed speed up (a hyperparameter that can be determined in a few quick trials) to make it harder to prune without loss in accuracy (Figure 8(a)), thereby exploiting the full potential of the metanetwork.
+This effect is analogous to removing low-quality samples from a dataset in order to obtain a higher-quality subset: an easy to prune network contains many low-quality parameters, which may harm the training of the metanetwork.
+This step isn’t necesary, but facilitates convergence, saves VRAM use, and helps achieve better pruning results (Figure 8(b)) during most of the time.
+(a) The origin network is easy to prune, after initial pruning and slight finetuning, it becomes hard to prune.
+(b) Initial pruning helps get better metanetworks and better pruning results
+Figure 8: With or without initial pruning (ResNet56 on CIFAR10 as an example)
+Metanetwork: Change the origin hard to prune network into a graph, feed it through the metanetwork to get a new graph, and change it back into a new network.
+Pruning: Prune the new easy to prune network.
+Finetuning: Train the network with a relatively small learning rate. This is a common step that is widely used in pruning tasks.
+Unlike prior learning to prune methods that require special training during each pruning, our pruning pipeline only requires a feed forward through metanetwork and standard finetuning, which is simple, general and effective.
+Once we have a metanetwork, we can prune as many networks as we want.
+All our pruning models are completely different from meta-training data models.
+This demonstrates our metanetwork naturally has great transferability and no prior work has done something like this before.
+Appendix D Experimental on CNNs
+D.1 General Experiment Setup
+D.1.1 General Settings
 Optimizer: For all data model training and fine-tuning, we use torch.optim.SGD with momentum=0.9.
@@
 To generate a data model, we train a base model, perform initial pruning followed by finetuning, and finally save the model along with relevant information.
-Typically, generating 3–10 data models is sufficient, with 2–8 used for meta-training and the remainder reserved for meta-evaluation or testing.
+Typically, generating 2–10 data models is sufficient, with 1–8 used for meta-training and the remainder reserved for meta-evaluation or testing.
 Meta Evaluation (Meta Eval): Meta evaluation is a process used to assess the quality of the metanetwork during meta-training.
@@
 Visualizing Metanetworks: This is a key technique used throughout our experiments.
-When we refer to visualizing a metanetwork, we mean passing a model through the metanetwork, followed by finetuning, and then visualize the “Test Accuracy vs.
+When we refer to visualizing a metanetwork, we mean passing a model through the metanetwork, followed by finetuning, and then visualize the “Acc VS.
 Speed-Up” curve of the resulting model, as illustrated in Figure 5.
@@
 For all other standard training and finetuning tasks, we use the small batch size.
-A.2 General Meta Training Details
+D.1.2 General Meta Training Details
 Equivalent Conversion between Network and Graph: During meta-training, we first convert the original network into an origin graph.
@@
 For untrainable parameters in the new network, we keep it the same as the old network.
-We provide more explanations for this conversion process in the following subsections: B, C, and D.
+We provide more explanations for this conversion process in the following subsections: D.2, D.3, and D.4.
 The Relative Importance of Sparsity Loss and Accuracy Loss: In our experiments, we introduce a hyperparameter called "pruner reg", which controls the relative importance of sparsity loss compared to accuracy loss.
 During backpropagation, the gradient from the sparsity loss is scaled(multiplied) by this "pruner reg" value, while the gradient from the accuracy loss remains unscaled.
-Meta Training Milestone: As discussed in the Ablation Study 3, as the number of training epochs increases, the metanetwork’s ability to make the network easier to prune becomes stronger, while its ability to maintain the accuracy becomes weaker.
+Meta Training Milestone: As discussed in the Ablation Study (Section 5.3), as the number of training epochs increases, the metanetwork’s ability to make the network easier to prune becomes stronger, while its ability to maintain the accuracy becomes weaker.
 During meta-training, we usually set a milestone for our learning rate scheduler.
-Before this milestone, we use a relatively large learning rate to quickly improve the metanetwork’s ability to make networks more prunable.
+Before this milestone, we use a relatively large learning rate to quickly improve the metanetwork’s ability to make networks easier to prune.
 After the milestone, we switch to a smaller learning rate to finetune the metanetwork and slow down the change of the two abilities, allowing us to select a well-balanced metanetwork.
@@
 We continue this process iteratively until we find a metanetwork that meets our criteria: a relatively high accuracy in the flat region and a long flat region indicating robustness to pruning.
-A.3 General Pruning Details :
-Pruning Speed Up: Once a suitable metanetwork has been selected for pruning, the next step is to determine the target speed up.
-In general, after visualizing the metanetwork, we observe a flat region followed by a sharp decline in accuracy.
-We define a turning point as the speed up at which the accuracy drops below a certain threshold.
-Experimentally, we find that we can safely prune the model at a speed up slightly higher than this turning point without incurring any accuracy loss after fine-tuning.
-For instance, when pruning ResNet-56 on CIFAR-10, we consider an accuracy drop below 0.93 as the turning point (typically around 2.0$\times$ speed up), and in practice, we are able to achieve a 2.9$\times$ speed-up with no loss in accuracy after finetuning.
-Appendix B ResNet56 on CIFAR10
-B.1 Equivalent Conversion between Network and Graph
+D.1.3 General Pruning Details :
+[…]
+D.2 ResNet56 on CIFAR10
+D.2.1 Equivalent Conversion between Network and Graph
 We change ResNet56 into a graph with node featrues of 8 dimensions and edge features of 9 dimensions.
@@
 If the residual connection includes parameters (e.g., a downsample with a convolutional layer and batch normalization), we construct the edge features in the same way as for standard convolutional layers.
-B.2 Meta Training
+D.2.2 Meta Training
 Data Models : We generate 10 models as our data models, among them, 8 are used for meta-training and 2 are used for validation (visualize the metanetwork).
@@
 Then we backward the gradients from both two losses to update our metanetwork.
-Training: We train our metanetwork with learning rate 0.001, milestone "12", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 28 as our final network for pruning.
-B.3 Full Results (Including Error Bars)
-All results are summarized in Table 4, where we repeat the pruning process 3 times using different seeds: 7, 8, 9.
-It is important to note that due to the design of our algorithm and implementation, we cannot prune the network to exactly the same speed up across different runs.
-For instance, when targeting a 1.3$\times$ speed up, the actual achieved speed up may be slightly larger, such as 1.31$\times$, 1.32$\times$, or 1.35$\times$.
+Training: We train our metanetwork with learning rate 0.001, milestone "3", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 39 as our final network for pruning.
+D.2.3 GPU Usage
+All tasks can be run on 1 NVIDIA RTX 4090.
+D.2.4 Full Results
+[…]
 The aggregated statistical results are presented in Table 7, where each value is reported in the form of mean(standard deviation).
+D.2.5 Hyperparameters
+See Table 8.
 Table 6: ResNet56 on CIFAR10 Full
 Method Base Pruned $\Delta$ Acc Pruned FLOPs Speed Up
-NISP [69] 43.2% 1.76
-Geometric [20] 93.59 93.26 -0.33 41.2% 1.70
-Polar [76] 93.80 93.83 0.03 46.8% 1.88
-DCP-Adapt [77] 93.80 93.81 0.01 47.0% 1.89
-CP [35] 92.80 91.80 -1.00 50.0% 2.00
-AMC [23] 92.80 91.90 -0.90 50.0% 2.00
-HRank [41] 93.26 92.17 -1.09 50.0% 2.00
-SFP [18] 93.59 93.36 -0.23 52.6% 2.11
-ResRep [8] 93.71 93.71 0.00 52.8% 2.12
-SCP [28] 93.69 93.23 -0.46 51.5% 2.06
-FPGM [19] 93.59 92.93 -0.66 52.6% 2.11
-FPC [21] 93.59 93.24 -0.35 52.9% 2.12
-DMC [11] 93.62 92.69 -0.93 50.0% 2.00
-GNN-RL [70] 93.49 93.59 0.10 54.0% 2.17
-DepGraph w/o SL [9] 93.53 93.46 -0.07 52.6% 2.11
-DepGraph with SL [9] 93.53 93.77 0.24 52.6% 2.11
-ATO [62] 93.50 93.74 0.24 55.0% 2.22
-Meta-Pruning (ours) 93.51 94.08 0.57 56.5% 2.30
-Meta-Pruning (ours) 93.51 93.56 0.05 56.5% 2.30
-Meta-Pruning (ours) 93.51 93.81 0.30 56.9% 2.32
-GBN [67] 93.10 92.77 -0.33 60.2% 2.51
-AFP [6] 93.93 92.94 -0.99 60.9% 2.56
-C-SGD [7] 93.39 93.44 0.05 60.8% 2.55
-Greg-1 [60] 93.36 93.18 -0.18 60.8% 2.55
-Greg-2 [60] 93.36 93.36 0.00 60.8% 2.55
-DepGraph w/o SL [9] 93.53 93.36 -0.17 60.2% 2.51
-DepGraph with SL [9] 93.53 93.64 0.11 61.1% 2.57
-ATO [62] 93.50 93.48 -0.02 65.3% 2.88
-Meta-pruning (ours) 93.51 93.69 0.18 65.6% 2.91
-Meta-pruning (ours) 93.51 93.42 -0.09 66.0% 2.94
-Meta-pruning (ours) 93.51 93.31 -0.20 66.6% 2.99
-Meta-pruning (ours) 93.51 93.25 -0.26 66.9% 3.02
-Meta-pruning (ours) 93.51 93.20 -0.31 66.7% 3.00
-Meta-pruning (ours) 93.51 93.18 -0.33 67.5% 3.08
+NISP (Yu et al., 2018) 43.2% 1.76
+Geometric (He et al., 2019b) 93.59 93.26 -0.33 41.2% 1.70
+Polar (Zhuang et al., 2020) 93.80 93.83 0.03 46.8% 1.88
+DCP-Adapt (Zhuang et al., 2018) 93.80 93.81 0.01 47.0% 1.89
+CP (Li et al., 2017a) 92.80 91.80 -1.00 50.0% 2.00
+AMC (He et al., 2018b) 92.80 91.90 -0.90 50.0% 2.00
+HRank (Lin et al., 2020a) 93.26 92.17 -1.09 50.0% 2.00
+SFP (He et al., 2018a) 93.59 93.36 -0.23 52.6% 2.11
+ResRep (Ding et al., 2021) 93.71 93.71 0.00 52.8% 2.12
+SCP (Kang & Han, 2020) 93.69 93.23 -0.46 51.5% 2.06
+FPGM (He et al., 2019a) 93.59 92.93 -0.66 52.6% 2.11
+FPC (He et al., 2020) 93.59 93.24 -0.35 52.9% 2.12
+DMC (Gao et al., 2020) 93.62 92.69 -0.93 50.0% 2.00
+GNN-RL (Yu et al., 2022) 93.49 93.59 0.10 54.0% 2.17
+DepGraph w/o SL (Fang et al., 2023) 93.53 93.46 -0.07 52.6% 2.11
+DepGraph with SL (Fang et al., 2023) 93.53 93.77 0.24 52.6% 2.11
+ATO (Wu et al., 2024) 93.50 93.74 0.24 55.0% 2.22
+Meta-Pruning (ours) 93.51 93.64 0.13 56.5% 2.30
+Meta-Pruning (ours) 93.51 93.75 0.24 58.0% 2.38
+Meta-Pruning (ours) 93.51 93.78 0.27 56.5% 2.30
+GBN (You et al., 2019a) 93.10 92.77 -0.33 60.2% 2.51
+AFP (Ding et al., 2018) 93.93 92.94 -0.99 60.9% 2.56
+C-SGD (Ding et al., 2019) 93.39 93.44 0.05 60.8% 2.55
+Greg-1 (Wang et al., 2021) 93.36 93.18 -0.18 60.8% 2.55
+Greg-2 (Wang et al., 2021) 93.36 93.36 0.00 60.8% 2.55
+DepGraph w/o SL (Fang et al., 2023) 93.53 93.36 -0.17 60.2% 2.51
+DepGraph with SL (Fang et al., 2023) 93.53 93.64 0.11 61.1% 2.57
+ATO (Wu et al., 2024) 93.50 93.48 -0.02 65.3% 2.88
+Meta-pruning (ours) 93.51 93.64 0.13 65.6% 2.91
+Meta-pruning (ours) 93.51 93.28 -0.23 65.9% 2.93
+Meta-pruning (ours) 93.51 93.49 -0.02 66.0% 2.94
+Meta-pruning (ours) 93.51 93.27 -0.24 66.8% 3.01
+Meta-pruning (ours) 93.51 93.10 -0.41 66.9% 3.02
+Meta-pruning (ours) 93.51 93.52 0.01 67.0% 3.03
 Table 7: ResNet56 on CIFAR10 Statistics
 Pruned Accuracy Pruned FLOPs Speed Up
-93.82($\pm$0.21) 56.63%($\pm$0.19%) 2.3067($\pm$0.0094)
-93.47($\pm$0.16) 66.07%($\pm$0.41%) 2.9467($\pm$0.0330)
-93.21($\pm$0.03) 67.03%($\pm$0.34%) 3.0333($\pm$0.0340)
-B.4 Computer Resources and Hyperparameters
-See Table 6.
+93.72($\pm$0.06) 57.0%($\pm$0.71%) 2.3267($\pm$0.0377)
+93.47($\pm$0.15) 65.83%($\pm$0.17%) 2.9267($\pm$0.0125)
+93.30($\pm$0.17) 66.90%($\pm$0.08%) 3.0200($\pm$0.0081)
 Table 8: ResNet56 on CIFAR10 Hyperparameters
@@
 weight decay 0.0005
-milestone "12"
+milestone "3"
 pruner reg 10
-Final Pruning metanetwork epoch 28
+Final Pruning metanetwork epoch 39
 speed up 2.3, 2.9, 3.0
@@
 milestone "80, 120"
-Appendix C VGG19 on CIFAR100
-C.1 Equivalent Conversion between Network and Graph
+D.3 VGG19 on CIFAR100
+D.3.1 Equivalent Conversion between Network and Graph
 We change VGG19 into a graph with node featrues of 5 dimensions and edge features of 9 dimensions.
@@
 The edge features are constructed by zero-padding all convolutional kernels to a size of $3\times 3$ (note that our networks only contain kernels of size $3\times 3$ or $1\times 1$), and then flattening them into feature vectors.
-C.2 Meta Training
+D.3.2 Meta Training
 Data Models : We generate 10 models as our data models, among them, 8 are used for meta-training and 2 are used for validation (visualize the metanetwork).
@@
 Then we backward the gradients from both two losses to update our metanetwork.
-Training: We train our metanetwork with learning rate 0.001, milestone "15", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 51 as our final network for pruning.
-C.3 Full Results (Including Error Bars)
-All results are summarized in Table 7, where we repeat the pruning process 3 times using different seeds: 7, 8, 9.
-It is important to note that due to the design of our algorithm and implementation, we cannot prune the network to exactly the same speed up across different runs.
-For instance, when targeting a 8.95$\times$ speed up, the actual achieved speed up may be slightly larger, such as 8.95$\times$, 9.01$\times$, or 9.02$\times$.
+Training: We train our metanetwork with learning rate 0.001, milestone "10", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 38 as our final network for pruning.
+D.3.3 GPU Usage
+All tasks can be run on 1 NVIDIA RTX 4090.
+D.3.4 Full Results
+[…]
 The aggregated statistical results are presented in Table 10, where each value is reported in the form of mean(standard deviation).
+D.3.5 Hyperparameters
+See Table 11.
 Table 9: VGG19 on CIFAR100 Full
 Method Base Pruned $\Delta$ Acc Pruned FLOPs Speed Up
-OBD [59] 73.34 60.70 -12.64 82.55% 5.73
-OBD [59] 73.34 60.66 -12.68 83.58% 6.09
-EigenD [59] 73.34 65.18 -8.16 88.64% 8.80
-Greg-1 [60] 74.02 67.55 -6.67 88.69% 8.84
-Greg-2 [60] 74.02 67.75 -6.27 88.69% 8.84
-DepGraph w/o SL [9] 73.50 67.60 -5.44 88.73% 8.87
-DepGraph with SL [9] 73.50 70.39 -3.11 88.79% 8.92
-Meta-Pruning (ours) 73.65 70.06 -3.59 88.85% 8.97
-Meta-Pruning (ours) 73.65 69.19 -4.46 88.91% 9.02
-Meta-Pruning (ours) 73.65 67.05 -6.60 89.07% 9.15
+OBD (Wang et al., 2019) 73.34 60.70 -12.64 82.55% 5.73
+OBD (Wang et al., 2019) 73.34 60.66 -12.68 83.58% 6.09
+EigenD (Wang et al., 2019) 73.34 65.18 -8.16 88.64% 8.80
+Greg-1 (Wang et al., 2021) 74.02 67.55 -6.67 88.69% 8.84
+Greg-2 (Wang et al., 2021) 74.02 67.75 -6.27 88.69% 8.84
+DepGraph w/o SL (Fang et al., 2023) 73.50 67.60 -5.44 88.73% 8.87
+DepGraph with SL (Fang et al., 2023) 73.50 70.39 -3.11 88.79% 8.92
+Meta-Pruning (ours) 73.65 68.65 -5.00 88.81% 8.94
+Meta-Pruning (ours) 73.65 67.63 -6.02 88.96% 9.06
+Meta-Pruning (ours) 73.65 69.75 -3.90 88.83% 8.95
 Table 10: VGG19 on CIFAR100 Statistics
 Pruned Accuracy Pruned FLOPs Speed Up
-68.77($\pm$1.26) 88.94%($\pm$0.09%) 9.0467($\pm$0.0759)
-C.4 Computer Resources and Hyperparameters
-See Table 9.
+68.68($\pm$0.87) 88.87%($\pm$0.07%) 8.9833($\pm$0.0544)
 Table 11: VGG19 on CIFAR100 Hyperparameters
@@
 in edge dim 9
-node res ratio 0.01
-edge res ratio 0.01
+node res ratio 0.05
+edge res ratio 0.05
 Meta Training lr 0.001
 weight decay 0.0005
-milestone "15"
+milestone "10"
 pruner reg 10
-Final Pruning metanetwork epoch 51
-speed up 8.95
+Final Pruning metanetwork epoch 38
+speed up 8.90
 Finetuning After Metanetwork epoch 2000
@@
 milestone "1850, 1950"
-Appendix D ResNet50 on ImageNet
-D.1 Equivalent Conversion between Network and Graph
+D.4 ResNet50 on ImageNet
+D.4.1 Equivalent Conversion between Network and Graph
 We change ResNet50 into a graph with node featrues of 8 dimensions and edge features of 1, 9 or 49 dimensions.
@@
 If the residual connection has no learnable parameters (i.e., it directly adds the input to the output), we represent the edge from $A_{i}$ to $B_{i}$ as a $1\times 1$ convolutional kernel with value 1.
-Different from section B, to reduce the VRAM footprint of the program, we build no edges between $A_{i}$ and $B_{j}$ ($i\neq j$).
+Different from section D.2, to reduce the VRAM footprint of the program, we build no edges between $A_{i}$ and $B_{j}$ ($i\neq j$).
 If the residual connection includes parameters (e.g., a downsample with a convolutional layer and batch normalization), we construct the edge features in the same way as for standard convolutional layers.
-D.2 Meta Training
+D.4.2 Meta Training
 Data Models : We generate 3 models as our data models, among them, 2 are used for meta-training and 1 are used for validation (visualize the metanetwork).
@@
 One Meta Training Epoch : We train with torch.nn.parallel.DistributedDataParallel (pytorch data parallel) across 8 gpus.
-Unlike in Section B and Section C, at the beginning of each epoch, we evenly distribute different data models across the GPUs.
+Unlike in Section D.2 and Section D.3, at the beginning of each epoch, we evenly distribute different data models across the GPUs.
 For example, since we have two data models, we load one on four GPUs and the other on the remaining four GPUs.
@@
 Each training epoch consists of (ImageNet data num / big batch size) iterations, meaning that one full pass over the entire ImageNet dataset when computing the accuracy loss is considered as one epoch .
-Training: We train our metanetwork with learning rate 0.01, milestone "4", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 13 as our final network for pruning.
-D.3 Full Results (Including Error Bars)
-All results are summarized in Table 10, where we repeat the pruning process 3 times.
-It is important to note that due to the design of our algorithm and implementation, we cannot prune the network to exactly the same speed up across different runs.
-For instance, when targeting a 2.3$\times$ speed up, the actual achieved speed up may be slightly larger, such as 2.31$\times$, 2.32$\times$, or 2.35$\times$.
+Training: We train our metanetwork with learning rate 0.01, milestone "2", weight decay 0.0005 and pruner reg 10. Finally we use metanetwork from epoch 12 as our final network for pruning.
+D.4.3 GPU Usage
+Meta-Training and feed forward through the metanetwork during pruning requires large VRAM and needs to be run on NVIDIA A100.
+All other training and finetuning can be run on NVIDIA RTX 4090.
+In practice, we use 8 gpus in parallel.
+D.4.4 Full Results
+[…]
 The aggregated statistical results are presented in Table 13, where each value is reported in the form of mean(standard deviation).
+D.4.5 Hyperparameters
+See Table 14.
 Table 12: ResNet50 on ImageNet Full
 Method Base Top-1(Top-5) Pruned Top-1($\Delta$) Pruned Top-5($\Delta$) Pruned FLOPs
-DCP [77] 76.01%(92.93%) 74.95%(-1.06%) 92.32%(-0.61%) 55.6%
-CCP [49] 76.15%(92.87%) 75.21%(-0.94%) 92.42%(-0.45%) 54.1%
-FPGM [19] 76.15%(92.87%) 74.83%(-1.32%) 92.32%(-0.55%) 53.5%
-ABCP [42] 76.01%(92.96%) 73.86%(-2.15%) 91.69%(-1.27%) 54.3%
-DMC [11] 76.15%(92.87%) 75.35%(-0.80%) 92.49%(-0.38%) 55.0%
-Random [38] 76.15%(92.87%) 75.13%(-1.02%) 92.52%(-0.35%) 51.0%
-DepGraph [9] 76.15%(-) 75.83%(-0.32%) - 51.7%
-ATO [62] 76.13%(92.86%) 76.59%(+0.46%) 93.24%(+0.38%) 55.2%
-DTP [39] 76.13%(-) 75.55%(-0.58%) - 56.7%
-ours 76.14%(93.11%) 76.17%(+0.03%) 92.94%(-0.17%) 56.9%
-ours 76.14%(93.11%) 76.27%(+0.13%) 93.05%(-0.06%) 56.9%
-ours 76.14%(93.11%) 76.15%(+0.01%) 92.95%(-0.16%) 56.9%
+DCP (Zhuang et al., 2018) 76.01%(92.93%) 74.95%(-1.06%) 92.32%(-0.61%) 55.6%
+CCP (Peng et al., 2019) 76.15%(92.87%) 75.21%(-0.94%) 92.42%(-0.45%) 54.1%
+FPGM (He et al., 2019a) 76.15%(92.87%) 74.83%(-1.32%) 92.32%(-0.55%) 53.5%
+ABCP (Lin et al., 2020b) 76.01%(92.96%) 73.86%(-2.15%) 91.69%(-1.27%) 54.3%
+DMC (Gao et al., 2020) 76.15%(92.87%) 75.35%(-0.80%) 92.49%(-0.38%) 55.0%
+Random (Li et al., 2022) 76.15%(92.87%) 75.13%(-1.02%) 92.52%(-0.35%) 51.0%
+DepGraph (Fang et al., 2023) 76.15%(-) 75.83%(-0.32%) - 51.7%
+ATO (Wu et al., 2024) 76.13%(92.86%) 76.59%(+0.46%) 93.24%(+0.38%) 55.2%
+DTP (Li et al., 2023) 76.13%(-) 75.55%(-0.58%) - 56.7%
+ours 76.14%(93.11%) 76.13%(-0.01%) 92.78%(-0.33%) 57.2%
+ours 76.14%(93.11%) 76.24%(+0.20%) 93.09%(-0.02%) 57.1%
+ours 76.14%(93.11%) 76.08%(-0.06%) 92.93%(-0.18%) 56.9%
 Table 13: ResNet50 on ImageNet Statistics
 Pruned FLOPs (Speed Up) Pruned Top-1 Pruned Top-5
-56.9%(2.32$\times$) 76.20%($\pm$0.05%) 92.98%($\pm$0.08%)
-D.4 Computer Resources and Hyperparameters
-See Table 12.
+[…]
 Table 14: ResNet50 on ImageNet Hyperparameters
 Types Name Value
-Compute Resources GPU NVIDIA A100 40G
+Compute Resources GPU NVIDIA A100 & NVIDIA RTX 4090
 parallel 8
@@
 milestone "30"
-Metanetwork num layer 4
-hiddim 11
+Metanetwork num layer 6
+hiddim 16
 in node dim 8
-node res ratio 0.001
-edge res ratio 0.001
+node res ratio 0.002
+edge res ratio 0.002
 Meta Training lr 0.01
 weight decay 0.0005
-milestone "4"
+milestone "2"
 pruner reg 10
-Final Pruning metanetwork epoch 13
+Final Pruning metanetwork epoch 12
 speed up 2.3095
@@
 Appendix E More About Ablation Study
-E.1 Visualize "Test Acc VS. Speed Up" Curves
-As mentioned in section 4.5, we visualize the "Test Acc VS. Speed Up" curves of metanetworks in Resnet56 on CIFAR10 in figure 4, VGG19 on CIFAR100 in figure 5 and ResNet 50 on ImageNet in figure 6.
-Figure 4: Resnet56 on CIFAR10 Visualization
-Figure 5: VGG19 on CIFAR100 Visualization
-Figure 6: ResNet50 on ImageNet Visualization
-E.2 Finetuning after Metanetwork
+E.1 Finetuning after Metanetwork
 (a) Finetune Different Epochs with Same Metanetwork
 (b) Use Finetune to Get Better Performance
-Figure 7: Finetune After Metanetwork (ResNet56 on CIFAR10)
+Figure 9: Finetune After Metanetwork (ResNet56 on CIFAR10 as example)
 As mentioned in section 5.1, our pruning pipeline can be summarized as:
-$InitialPruning(Finetuning)\rightarrow Metanetwork(Finetuning)\rightarrow Pruning(Finetuning)$
+$\underbrace{\text{Initial Pruning (Finetuning)}}_{\text{Optional}}\rightarrow\underbrace{\text{Metanetwork (Finetuning)}\rightarrow\text{Pruning (Finetuning)}}_{\text{Necessary}}$ (29)
 In this pipeline, finetuning after pruning is a common way to improve the accuracy.
@@
 To illustrate this, we use the example of pruning ResNet56 on CIFAR10.
-In Figure 7(a), we visualize the "Test Accuracy vs.
+In Figure 9(a), we visualize the "Acc vs.
 Speed Up" curves for models after the same metanetwork and after various amounts of finetuning.
@@
 our accuracy loss in meta-training works)
-  • Trend of Change: As the number of finetuning epochs increases, the flat portion of the curve becomes shorter, and the overall accuracy of this region improves.
+  • Trend of Change: As the number of finetuning epochs increases, the flat portion of the curve becomes shorter or stays the same, and the overall accuracy of this region improves.
 When we feed the same network into metanetworks trained for different numbers of meta-training epochs and get several modified networks, then finetune them with different epochs to reach the same accuracy (as shown in Figure 9(b)), another important pattern emerges:
@@
 This is evident in Figure 9(b), where the green line (representing a higher meta-training epoch) exhibits a longer flat region, indicating improved robustness to pruning.
-This suggests that the effectiveness of our method can be further improved by employing metanetworks trained for more meta-training epochs and applying additional finetuning afterward.
+This suggests that the effectiveness of our method can be further improved by employing metanetworks trained for more meta-training epochs and applying more finetuning afterward.
 In practice, however, we avoid excessive finetuning in order to keep the number of finetuning epochs within a reasonable range, allowing for multiple experimental trials.
 Overall, all finetuning in our pipeline is necessary and effective.
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The Abstract and Introduction sections accurately reflect the contribution and scope of the paper.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: The limitations of our work is discussed in the conclusion section (5).
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [N/A]
-  14. Justification: We propose no theoretical result that needs to be proved.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: We open sourced our code and provide everything you need to reproduce our code in Appendix.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: We provides a URL link to our released code. All information about datasets and how to run our code can be found together with our code.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: In Ablation Study and Appendix we talk about all the details of our experiments and how to choose hyparameter.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: We run each experiment 3 times and calculate the error bars, which you can see in Appendix.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: We report our computer resources for each experiment in Appendix
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: This paper conforms to the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A]
-  49. Justification: There is no societal impact of the work.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: This paper has no risks and needs no safeguards.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: The assets utilized in this paper are appropriately acknowledged. The associated license and terms of use have been clearly stated and duly adhered to.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [Yes]
-  64. Justification: We open sourced our code and model with clear explanations.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: We involve no crowdsourcing subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: We involve no crowdsourcing subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: We didn’t involve LLMs as any important, original, or non-standard components.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Appendix F Experiments on Transferability
+F.1 Transfer between datasets
+The common pruning pipeline is:
+$\text{Initial Pruning (Finetuning)}\rightarrow\text{Metanetwork (Finetuning)}\rightarrow\text{Pruning (Finetuning)}$ (30)
+Pruning pipeline for None is:
+Initial $Pruning (Finetuning)\text${Initial $\rightarrow\text{Pruning (Finetuning)}$ (31)
+The dataset of a to be pruned network means we do all training and finetuning on it using this dataset.
+The dataset of a metanetwork means we generate dataset models, meta-train the metanetwork using this dataset.
+We mainly have 2 hyperparameters here–initial speed up and final speed up .
+We approximately set them based on the difficulty of each dataset.
+Here dataset refers to the dataset of the to be pruned network (rows in Table 15).
+For CIFAR10, the initial pruning speed up is 1.32x and the final speed up is 3.0x.
+For CIFAR100 the initial pruning speed up is 1.32x and the final speed up is 2.5x.
+For SVHN, the initial pruning speed up is 3.0x and the final speed up is 10.0x.
+Table 15: Transfer between datasets: All networks’ architecture is ResNet56.
+Columns represent the training datasets for the metanetwork, and rows represent the training datasets for the to be pruned network.
+“None” indicates using no metanetwork.
+Results with metanetwork is obviously better than no metanetwork (The only exception is when training datasets for the to be pruned network is SVHN, and we guess this is because the dataset SVHN itself is too easy).
+[…]
+CIFAR10 93.35 92.47 92.87 91.28
+CIFAR100 69.97 70.16 69.25 68.91
+SVHN 96.79 96.50 96.86 96.78
+F.2 Transfer between Architectures
+The common pruning pipeline is:
+$\text{Initial Pruning (Finetuning)}\rightarrow\text{Metanetwork (Finetuning)}\rightarrow\text{Pruning (Finetuning)}$ (32)
+Pruning pipeline for None is:
+Initial $Pruning (Finetuning)\text${Initial $\rightarrow\text{Pruning (Finetuning)}$ (33)
+The architecture of a to be pruned network means this network is constructed using this architecture.
+The architecture of a metanetwork means all dataset models we generated for meta-training this metanetwork use this architecture.
+We mainly have 2 hyperparameters here–initial speed up and final speed up.
+We approximately set them based on the ability of each architecture.
+Here architecture refers to the architecture of the to be pruned network (rows in Table 16).
+For ResNet56, the initial pruning speed up is 1.32x and the final speed up is 3.0x.
+For ResNet110 the initial pruning speed up is 2.0x and the final speed up is 4.0x.
+Table 16: Transfer between architectures.
+All training dataset is CIFAR10.
+Columns represent the architectures used for training the metanetwork, and rows represent the architecures of the to be pruned network.
+“None” indicates using no metanetwork.
+All results with metanetwork is obviously better than no metanetwork.
+[…]
+ResNet56 93.40 92.81 92.08
+ResNet110 93.04 93.38 92.40
+Appendix G Expriments on Transformers
+G.1 MHSA to Graph
+[…]
+$Q_{h}=\mathbf{X}\mathbf{W}_{h}^{Q},\quad K_{h}=\mathbf{X}\mathbf{W}_{h}^{K},\quad V_{h}=\mathbf{X}\mathbf{W}_{h}^{V}.$
+Then each head computes attention via
+$Y_{h}=\text{softmax}(Q_{h}K_{h}^{\top})V_{h},$
+The outputs $Y_{1},\ldots,Y_{H}$ are concatenated and linearly projected:
+$\text{MHSA}(\mathbf{X})=\text{Concat}(Y_{1},\ldots,Y_{H})\mathbf{W}^{O}.$
+The input and output are both $d$-dimensional.
+Each head produces $d_{H}$-dimensional outputs.
+When changing to a graph, we represent this with $d$ input nodes in the first layer, $H\cdot d_{H}$ attention head nodes in the second layer, and $d$ output nodes in the third layer.
+Between the first and the second layer, we model the three projection types (query, key, value) using multidimensional edge features: for each edge $(i,j)$ in head $h$, the feature is
+$e_{ij}^{h}=\big((\mathbf{W}_{h}^{Q})_{ij},\;(\mathbf{W}_{h}^{K})_{ij},\;(\mathbf{W}_{h}^{V})_{ij}\big).$
+Between the second and third layer, concatenation and the final projection $\mathbf{W}^{O}$ are handled naturally by the graph structure and treated as a standard linear layer.
+In summary, every pair of nodes between the first and second layers is connected by an edge characterized by three-dimensional features, which represent the attention parameters.
+Each pair of nodes between the second and third layers is connected by an edge with a single-dimensional feature, corresponding to the final linear projection.
+G.2 Equivalent Conversion between Network and Graph
+We change ViT-B/16 into a graph with node featrues of 6 dimensions and edge features of 1, 3 or 256 dimensions.
+We employ 3 linear layers to project edge features of these 3 different dimensions into the same hidden dimension, enabling their integration into our graph neural network.
+The node features consist of 6 features, they are weight and bias of previous layer norm, bias of previous linear layer, biases of query, key, value of previous attention layer.
+In the case that any of these features doesn’t exist, they are replaced by a default value of $[1,0,0,0,0,0]$ at their corresponding positions.
+The edge features are constructed almost in the same way as previous experiments. The only new part is MHSA, and we construct it as described in Appendix G.1.
+G.3 Meta Training
+Data Models: We use the default ViT-B/16 provided by PyTorch as the only data model.
+All training and pruning are conducted on this model, as training several separate models from scratch would be prohibitively time-consuming.
+Moreover, this choice ensures consistency with prior work in the field.
+Based on the Acc VS.
+speed up curve of the original ViT (Figure 10), we do initial pruning with a speed up of 1.0370x and apply no finetuning.
+This configuration yields a network with an accuracy of 81%.
+Figure 10: Acc VS. Speed Up curve of the origin ViT
+One Meta Training Epoch : The same as ResNet50 on ImageNet (Appendix D.4.2).
+Training: We train our metanetwork with learning rate 0.01, milestone "3", weight decay 0.0005 and pruner reg 1000. Finally we use metanetwork from epoch 6 as our final network for pruning.
+G.4 GPU Usage
+Meta-Training and feed forward through the metanetwork during pruning requires large VRAM and needs to be run on NVIDIA A100.
+All other training and finetuning can be run on NVIDIA RTX 4090.
+In practice, we use 8 gpus in parallel.
+G.5 Full Results
+All results are summarized in Table 17.
+Note that unlike CNNs, Vision Transformers naturally lacks inductive bias and need much more epochs of finetuning to reach higher accuracy.
+For other works like DepGraph (Fang et al., 2023), theirs method need to finetune 600 epochs in total, among them, 300 epochs are sparsity learning (slower than usual finetuning) and it takes more than a week to run a single round.
+We don’t have enough time to fully run the experiments with more finetuning for now, but we will continue tune our hyperparameters, add more epochs and hopefully give a better results in the next version of our paper.
+G.6 Hyperparameters
+See Table 18.
+Table 17: ViT-B/16 on ImageNet
+[…]
+ViT-B/16 (Dosovitskiy et al., 2021) 81.07% - - 17.6
+CP-ViT (Song et al., 2022) 77.91% 77.36% -0.55% 11.7
+DepGraph (Fang et al., 2023) 81.07% 79.17% -1.90% 10.4
+MetaPruning(ours) 81.07% 77.59% -2.83% 11.0
+Table 18: ViT-B/16 on ImageNet Hyperparameters
+Types Name Value
+Compute Resources GPU NVIDIA A100 & NVIDIA RTX 4090
+parallel 8
+Batch size small batch size 128 $\times$ 8
+big batch size 32 $\times$ 8 $\times$ 200
+Prepare Data Models data model num 1
+Generate Data Model epoch 0
+lr -
+weight decay -
+milestone -
+Initial Pruning speed up 1.0370
+Finetuning epoch 0
+lr -
+weight decay -
+milestone -
+Metanetwork num layer 3
+hiddim 4
+in node dim 6
+node res ratio 0.1
+edge res ratio 0.1
+Meta Training lr 0.01
+weight decay 0.0005
+milestone "3"
+pruner reg 1000
+Final Pruning metanetwork epoch 6
+speed up 1.6000
+Finetuning After Metanetwork epoch 100
+lr 0.0001
+weight decay 0.01
+scheduler cosineannealinglr
+Finetuning After Pruning epoch 100
+lr 0.0001
+weight decay 0.0001
+scheduler cosineannealinglr
+Appendix H The Use of Large Language Models
+We only use LLMs to check grammar errors and polish our writing. All other works are done by human writers.

### v3

@@
 1 Introduction
-With the rapid advancement of deep learning (LeCun et al., 2015; Schmidhuber, 2015), neural networks have become increasingly powerful.
+With the rapid advancement of deep learning (DeepLearning; DeepLearningOverview), neural networks have become increasingly powerful.
 However, this improved performance often comes with a significant increase in the number of parameters and computational cost (FLOPs).
 As a result, there is growing interest in methods to simplify these networks while preserving their performance.
-Pruning, which involves selectively removing certain parts of a neural network, has proven to be an effective approach (Cheng et al., 2024; He & Xiao, 2024; Reed, 1993).
+Pruning, which involves selectively removing certain parts of a neural network, has proven to be an effective approach (PruningSurvey; PruningSurveyStructured; PruningSurveyOld).
 A key idea of a large number of previous pruning works is to remove the unimportant components of a neural network.
 To achieve this goal, various criteria that measure the importance of components of a neural network have been invented.
-Some criteria are based on norm or magnitude scores (Han et al., 2016; Li et al., 2017b; Sun et al., 2024; He et al., 2018a).
-Some are based on how much the performance changes when certain parts of the network are removed or preserved (You et al., 2019b; Ma et al., 2023; Frankle & Carbin, 2019; Ye et al., 2020).
-Some others defined more complex criteria such as saliency and sensitivity (LeCun et al., 1989; Lee et al., 2019; Zhao et al., 2019).
+Some criteria are based on norm or magnitude scores (DeepCompression; PruningFiltersForEfficientConvNets; NormLLM; SFP).
+Some are based on how much the performance changes when certain parts of the network are removed or preserved (GateDecorator; LLMPruner; LotteryTicket; SubnetworkPruning).
+Some others defined more complex criteria such as saliency and sensitivity (OptimalBrainDemage; Snip; VariationalConvolutionalNeuralNetworkPruning).
 Once a pruning criterion is established, we can assess whether a model is easy or hard to prune based on that criterion.
@@
 However, several studies adopting this latter perspective have achieved promising results.
-Notably, sparsity regularization methods (Wen et al., 2016; Fang et al., 2023; Wang et al., 2021) exemplify this approach by encouraging models to become sparser, thereby facilitating more effective pruning.
+Notably, sparsity regularization methods (LearningStructuredSparsityInDeepNeuralNetworks; DepGraph; Greg) exemplify this approach by encouraging models to become sparser, thereby facilitating more effective pruning.
 Our framework inherits the idea of the second way.
@@
 Using neural networks to process other neural networks has emerged as an intriguing research direction.
-Early studies demonstrated that neural networks can extract useful information directly from the weights of other networks (Unterthiner et al., 2020).
+Early studies demonstrated that neural networks can extract useful information directly from the weights of other networks (AccFromWeights).
 We define a metanetwork as a neural network that takes another neural network as input and outputs either information about it or a modified network.
-Initial metanetworks were simple multilayer perceptrons (MLPs), while more recent designs incorporate stronger inductive biases by preserving symmetries (Godfrey et al., 2022) inherent in the input networks.
+Initial metanetworks were simple multilayer perceptrons (MLPs), while more recent designs incorporate stronger inductive biases by preserving symmetries (SymmetryDeepLearning) inherent in the input networks.
 Broadly, metanetwork architectures can be categorized into two perspectives.
-The weight space view applies specially designed MLPs directly on the model’s weights (Zaheer et al., 2017; Navon et al., 2023; Zhou et al., 2023a; Tran et al., 2024b; Zhou et al., 2023b; 2024; Tran et al., 2024a).
-And the graph view transforms the input network into a graph and applies graph neural networks (GNNs) to it (Lim et al., 2024; Kofinas et al., 2024; Kalogeropoulos et al., 2024).
-In this work, we adopt the graph view to build our metanetwork, leveraging the natural correspondence between graph nodes and neurons in a network layer, as well as the symmetries they exhibit (Maron et al., 2019).
+The weight space view applies specially designed MLPs directly on the model’s weights (DeepSets; DWS; NFN; NFNTransformer; NFTransformer; UNF; MonomialNFN).
+And the graph view transforms the input network into a graph and applies graph neural networks (GNNs) to it (GMN; NGGNN; ScaleGMN).
+In this work, we adopt the graph view to build our metanetwork, leveraging the natural correspondence between graph nodes and neurons in a network layer, as well as the symmetries they exhibit (InvariantEquivariantGNN).
 Given the maturity and effectiveness of GNNs, the graph metanetwork approach offers both elegance and practicality.
 2.2 Graph Neural Networks
-Graph neural networks (GNNs) are a class of neural networks specifically designed to operate on graph-structured data, where graphs consist of nodes and edges with associated features (Wu et al., 2021; Scarselli et al., 2009; Kipf & Welling, 2017).
-GNNs have gained significant attention in recent years, with frameworks such as the message passing neural network (MPNN) (Gilmer et al., 2017) proving to be highly effective.
-In this paper, we employ the message passing framework, specifically using Principal Neighborhood Aggregation (PNA) (Corso et al., 2020) as the backbone architecture for our metanetwork.
+Graph neural networks (GNNs) are a class of neural networks specifically designed to operate on graph-structured data, where graphs consist of nodes and edges with associated features (GNNSurvey; TheGraphNeuralNetworkModel; SemiSupervisedClassificationWithGraphConvolutionalNetworks).
+GNNs have gained significant attention in recent years, with frameworks such as the message passing neural network (MPNN) (MessagePassing) proving to be highly effective.
+In this paper, we employ the message passing framework, specifically using Principal Neighborhood Aggregation (PNA) (PNA) as the backbone architecture for our metanetwork.
 2.3 Sparsity Regularization Based Pruning
@@
 Prior research has explored various regularization techniques to encourage sparsity in neural networks to facilitate pruning.
-This idea of sparsity regularization has been widely adopted across many works (Wen et al., 2016; Gordon et al., 2018; Huang & Wang, 2018; Lin et al., 2020c; He et al., 2017; Lin et al., 2019; Xia et al., 2022; 2024; Fang et al., 2023; Wang et al., 2021).
+This idea of sparsity regularization has been widely adopted across many works (LearningStructuredSparsityInDeepNeuralNetworks; MorphNet; DataDrivenSparseStructureSelection; TowardCompactConvNetsViaStructureSparsityRegularizedFilterPruning; ChannelPruningForAcceleratingVeryDeepNeuralNetworks; TowardsOptimalStructuredCNNPruningViaGenerativeAdversarial; StructuredPruningLearnsCompactAndAccurateModels; ShearedLLaMA; DepGraph; Greg).
 Our method inherits this principle to some extent, as our metanetwork can be viewed as transforming the original network into a sparser version for better pruning.
 2.4 Learning to Prune & Meta Pruning
-Given the complexity of pruning, several previous works have leveraged neural networks (Dery et al., 2024; He et al., 2017; Wu et al., 2024; Chen et al., 2023), reinforcement learning (Rao et al., 2019; Yu et al., 2022; He et al., 2018b), and other techniques to automatically learn pruning strategies.
-Viewing pruning as a learning problem naturally leads to the idea of meta-learning, which learns “how to learn” (Hospedales et al., 2022).
-Some prior works have specifically applied meta-learning to pruning (Liu et al., 2019; Li et al., 2020).
+Given the complexity of pruning, several previous works have leveraged neural networks (EverybodyPruneNow; ChannelPruningForAcceleratingVeryDeepNeuralNetworks; ATO; OTO), reinforcement learning (RuntimeNetworkRoutingForEfficientImageClassification; GNN-RL; AMC), and other techniques to automatically learn pruning strategies.
+Viewing pruning as a learning problem naturally leads to the idea of meta-learning, which learns “how to learn” (MetaLearningSurvey).
+Some prior works have specifically applied meta-learning to pruning (MetaPruningFormer; DHP).
 Our work also draws idea from meta-learning, but it learns in a way entirely different from all prior works.
@@
 Conversion between networks and graphs: To apply a metanetwork (GNN) on networks, we first need to establish a conversion between networks and graphs.
-Prior works such as Kofinas et al.
-(2024); Lim et al.
-(2024) have managed to change networks of almost all architectures (CNN, Transformer, RNN, etc.) into graphs.
-Here we show how we establish conversion between a ResNet (He et al., 2016) and a graph in our experiments as an example.
-This inherits the idea of Kofinas et al.
-(2024); Lim et al.
-(2024) but is quite different in implementation details.
+Prior works such as NGGNN; GMN have managed to change networks of almost all architectures (CNN, Transformer, RNN, etc.) into graphs.
+Here we show how we establish conversion between a ResNet (Resnet) and a graph in our experiments as an example.
+This inherits the idea of NGGNN; GMN but is quite different in implementation details.
 To convert a network into a graph, we establish the correspondence between the components of a neural network and elements of a graph as follows:
@@
 This includes fully connected layers, convolutional layers, and residual connections.
-(3) Node Features: Node features comprise parameters associated with neurons, here we use the weight, bias, running mean, running var of batchnorm (Ioffe & Szegedy, 2015), including both batchnorms from previous adjacent layers and previous skip connection layers (if exist).
+(3) Node Features: Node features comprise parameters associated with neurons, here we use the weight, bias, running mean, running var of batchnorm (BatchNorm), including both batchnorms from previous adjacent layers and previous skip connection layers (if exist).
 (4) Edge Features: Edge features encode the weights of connections between nodes.
@@
 Our metanetwork must be powerful enough to learn the transforming rules for better pruning.
-We build our metanetwork based on the message passing framework (Gilmer et al., 2017) and PNA (Corso et al., 2020) architecture.
+We build our metanetwork based on the message passing framework (MessagePassing) and PNA (PNA) architecture.
 A brief overview of our architecture is provided below, the full description is available in Appendix B.1.
@@
 Optimizing the network using this loss can make it easier to prune.
-We draw ideas from prior sparsity regularization based pruning methods(section 2.3), and design our pruning criterion based on them (especially Fang et al.
-(2023)).
+We perform most of our experiments by default using structural pruning, but also include a small number of experiments with unstructured and N:M sparsity pruning to demonstrate that our method can be applied to all kinds of pruning.
+In unstructured pruning and N:M sparisity pruning, we simply use the naive $l_{1}$ norm as both our importance score function and sparsity loss.
+In structural pruning, we draw ideas from prior sparsity regularization based pruning methods(section 2.3), and design our pruning criterion based on them (especially DepGraph).
 We will give a brief introduction below.
 For a rigorous mathematical definition and more details, please refer to appendix B.2.
-We only do structural pruning in all our experiments.
 A variant of the classical $\ell_{2}$ norm, the group norm is used as our importance score function.
@@
 The choice of pruning criteria is highly flexible (see in our later experiments, section 6.1), which further demonstrates the generality of our framework.
-Although we conduct only structural pruning in all our experiments, all our methods can be directly applied to unstructured pruning as well.
+We perform most of our experiments by default using structural pruning, but also include a small number of experiments with unstructured and N:M sparsity pruning to demonstrate that our method can be applied to all kinds of pruning.
 4.4 A Holistic Perspective
@@
 Universally applicability: Theoretically, our implementations can be applied to almost all types of networks with all kinds of pruning.
-Almost all types of networks can be converted into graphs (Kofinas et al., 2024; Lim et al., 2024), and the designs of the GNN and the pruning criteria are also broadly compatible.
+Almost all types of networks can be converted into graphs (NGGNN; GMN), and the designs of the GNN and the pruning criteria are also broadly compatible.
 This makes our framework a universally applicable solution for network pruning.
@@
 This demonstrates our metanetwork naturally has great transferability and no prior work has done something like this before.
-See Appendix A for more comparisons between our work and prior ones.
+See Appendix A for more comparisons between our work and prior ones (ideas, costs etc.).
+In all experiments, we adopt structured pruning by default. Unstructured pruning will be explicitly stated when used (Section 6.2).
 5.2 Classical CNN Pruning Tasks
-We carry out our experiments on three most classical, popular, and representative image recognition tasks, including pruning ResNet56 (He et al., 2016) on CIFAR10 (Krizhevsky & Hinton, 2009), VGG19 (Simonyan & Zisserman, 2015) on CIFAR100 (Krizhevsky & Hinton, 2009) and ResNet50 (He et al., 2016) on ImageNet (Deng et al., 2009).
+We carry out our experiments on three most classical, popular, and representative image recognition tasks, including pruning ResNet56 (Resnet) on CIFAR10 (CIFAR), VGG19 (VGG) on CIFAR100 (CIFAR) and ResNet50 (Resnet) on ImageNet (IMAGENET).
 See appendix D.1 for more general setups.
@@
 (3) 76.14%(93.11%) 76.13%(-0.01%) 92.78%(-0.33%) 57.2%
-5.3 Ablation Study
+5.3 Ablation Study of General Behavioral Tendencies
 How metanetwork works: We visualized the "Acc vs.
@@
 Finetuning after Metanetwork. See appendix E.1 for how number of finetuning epochs after metanetwork influence the pruning results.
+5.4 Ablation Study of Statistics
+We compare the statistics between the origin network and the network after metanetwork(finetuning) to find out how metanetwork transforms the network to make it easier to prune.
+We mainly visualize the $l2$ norm and taylor sensitivity distribution of each layer in Figure 7.
+Taylor sensitivity here is defined as $w\cdot\Delta L$, it estimates how much the loss will increase if we mask this weight to zero.
+The larger taylor sensitivity, the more important the weight is and we are unlikely to prune it.
+More visualization of other statistics are in Appendix H
+Figure 7: Statistics: We compare the statistics of the origin network (orange) and the network after metanetwork and finetuning (blue) when pruning ResNet56 on CIFAR10.
+The first row is layerwise mean $l_{2}$ norm and taylor sensitivity.
+Then we randomly select several layers and visualize their $l_{2}$ norm distribution in row 2 and taylor sensitivity distribution in row 3.
+Observation: Row 1 shows both norm and sensitivity drops on average, especially in the latter part of network where contains a large amount of redundancy and can be substantially pruned.
+Row 2 shows $l_{2}$ norm distribution has been greatly changed.
+Large norms still exist but more norms tend to be very small.
+Row 3 shows more parameters become less important under taylor sensitivity.
 6 Experiments on Transferability and Flexibility
@@
 We build a series of reasonable pruning criterion (Appendix B.2).
-At first, we directly use MEAN REDUCE, MAX NORMALIZE, $\alpha=4$ as default, because this is the same as group norm in Fang et al.
-(2023).
+At first, we directly use MEAN REDUCE, MAX NORMALIZE, $\alpha=4$ as default, because this is the same as group norm in DepGraph.
 Later, we keep everything else the same, and try many different criteria, and find they all works well (Table 2).
@@
 Naive NONE 0 93.04% 65.75% 2.92
+6.2 Unstructured Pruning & N:M Sparsity Pruning
+Most of our experiments are conducted using structured pruning as default.
+Here we conduct unstructured pruning and N:M sparsity pruning to demonstrate that our methods can be use in all kinds of pruning.
+We don’t use FLOPs to measure the pruning results because unstructured pruning reduces no FLOPs and needs specialized algorithm to accelerate, so we measure the number of parameters instead.
+See results in Table 3
+Table 3: Unstructured & N:M sparsity pruning: Pruning ResNet56 on CIFAR10, meta-train and prune directly with the classical $l_{1}$ norm.
+From the results we can see (1) our framework also works great on unstructured pruning.
+(2) Unstructured pruning outperforms structured pruning because it is more flexible.
+Methods Pruned Acc Left Params Methods Pruned Acc Left Params
+Unstructured 93.95% 50.25% Structured(2.9x) 93.49% 42,96%
+94.14% 20.40% 3:4 94.02% 75.14%
+93.43% 15.42% 2:4 93.97% 50.27%
+92.96% 10.45% 1:4 93.13% 25.41%
 6.3 Transfer between Datasets and Architectures
-Section 4.4 mentioned that our metanetwork has natural transferability. Following experiments demonstrate that our metanetwork can transfer between similar datasets and network architectures.
-Transfer between datasets: 3 datasets, CIFAR10 (Krizhevsky & Hinton, 2009), CIFAR100 (Krizhevsky & Hinton, 2009) and SVHN (Goodfellow et al., 2014) are used.
+Section 4.4 mentioned that our metanetwork has natural transferability.
+Following experiments demonstrate that our metanetwork can transfer between similar datasets and network architectures.
+More relative experiments are in Appendix F.3.
+Transfer between datasets: 3 datasets, CIFAR10 (CIFAR), CIFAR100 (CIFAR) and SVHN (SVHN) are used.
 We train metanetwork and the to be pruned network on each of them, traverse through every possible combination, and use the metanetwork to prune the network.
@@
 SVHN 96.79 96.50 96.86 96.78
-Transfer between architectures: 2 architectures, ResNet56 and ResNet110 (He et al., 2016) are used.
+Transfer between architectures: 2 architectures, ResNet56 and ResNet110 (Resnet) are used.
 We train metanetwork and the to be pruned network on each of them, traverse through every possible combination, and use the metanetwork to prune the network.
@@
 As mentioned in section 4.4, our framework is theoretically applicable to almost all types of networks with all kinds of pruning.
-Here we expand our implementations to another widely used arthitecture–transformer (Vaswani et al., 2017).
-More specifically, we pruned the vision transformer (ViT) from Dosovitskiy et al.
-(2021) that is trained on ImageNet.
+Here we expand our implementations to another widely used arthitecture–transformer (Transformer).
+More specifically, we pruned the vision transformer (ViT) from ViT that is trained on ImageNet.
 See Appendix G for full experiments, including how we convert transformer into graph and conduct further meta-training and pruning.
@@
 One important question for almost all previous meta-learning approaches is the computational and memory costs are too large.
-We provide some estimate of memory and computational costs for scaling our methods to larger models and compare with two classic pruning methods–pruning at initalization methods (cheap pruning, hard to training and low accuracy like Lee et al.
-(2019); Wang et al.
-(2022)) and iterative magnitude pruning (costly, pruning during training and high performance like Frankle & Carbin (2019); Molchanov et al.
-(2019)).
+We provide some estimate of memory and computational costs for scaling our methods to larger models and compare with two classic pruning methods–pruning at initalization methods (cheap pruning, hard to training and low accuracy like Snip; pruningatinitsurvey) and iterative magnitude pruning (costly, pruning during training and high performance like LotteryTicket; ImportanceEstimationforNNP).
 The results show that our method achieves outstanding results while uses relatively low costs.
@@
 A.3 Generality
-Almost all previous meta-learning based pruning approaches, such as Liu et al.
-(2019); Li et al.
-(2020); Wu et al.
-(2024), are tailored to a specific network and thus lack generality.
+Almost all previous meta-learning based pruning approaches, such as MetaPruningFormer; DHP; ATO, are tailored to a specific network and thus lack generality.
 Because of this they require specialized training for each pruning instance.
@@
 Once our metanetwork is trained, it can be used to prune as many networks as we want, and even transfer between datasets and architectures.
+We divided the generality of learning to prune methods into 4 stages.
+  1. (1)
+    Learning once prune one specific network.
+  2. (2)
+    Learning once prune one type of networks(same architecture and dataset) as many as we want.
+  3. (3)
+    Learning once prune one group of networks(similar architectures and datasets) as many as we want.
+  4. (4)
+    Learning once prune any networks.
+Here learning refers to learn with extra trainings like gradient descent, reinforcement learning, etc.
+Finetuning or other fixed rules processes are not included.
+As far as we know, all prior learning to prune methods only reach stage (1).
+Our method reaches stage (3) and shows great improvement in generality over prior works.
+Follow our ideas and pretrain the metanetwork on various networks and datasets may provide a possible way to state (4), whether this will work requires further exploration in the future.
+In summary, our method shows great improvement in generality over prior learning to prune methods.
+A.4 A Concrete Example
+We provide a concrete example compared with prior works and report everything-time, hardware, gpu memory, results, etc. to give readers a more intuitive understanding of our method.
+We compare our work with DepGraph on pruning ResNet56 on CIFAR10.
+We choose DepGraph because we want to compare our work with learning to prune methods in recent years that also have strong results like us.
+The best candidates are DepGraph and ATO.
+While ATO is a pruning before training method, both DepGraph and our work are pruning after training methods, so we choose DepGraph.
+The key idea of DepGraph is sparsity training before pruning.
+It does special training on the network before pruning to make it more sparse and easier to prune.
+All experiments are run on 1 NVIDIA RTX 4090.
+See table 7 for time consumptions of dfferent methods.
+To align with DepGraph, we don’t use init pruning, and target at a speed up of 2.5x, which is the largest speed up in paper DepGraph.
+In ours(full), we finetune 100 epochs after metanetwork and 100 epochs after pruning.
+But later we found our method is stronger and the speed up 2.5x isn’t that hard.
+So in ours(efficient), we finetune 60 epochs after metanetwork and 60 epochs after pruning but also get results comparable to DepGraph.
+The DepGraph experiments use their default settings, sparsity training 100 epochs before pruning and finetune 100 epochs after pruning.
+Table 7: Time (miniutes)
+Process Ours(full) Ours(efficient) DepGraph
+Meta-Train 192 + 165 = 357 192 + 165 = 357 0
+Prune & Finetune 67 43 84
+From the results we can see our methods is quite efficient.
+It uses 357 minutes for meta-train, which is in a resonable range.
+We can greatly reduce time consumption in this process with some experiences, but for the sake of fairness in comparison, we must pretend we have no relevant experience.
+Among the 357 minutes, 192 minutes are used for generating 100 epochs metanetworks and 165 minutes are used for visualizing the Acc VS.
+Speed Up Curve to select the appropriate metanetwork for pruning.
+Once a metanetwork is trained, it can be used to prune as many networks as we want, so we can amortize the time for meta-training into each pruning.
+Figure 8 shows the amortized time consumption of our work compared to DepGraph.
+In all, our work is more powerful and more efficient in time if we prune several more networks or already have a trained metanetwork.
+Figure 8: Amortized Time Consumption
+As we mentioned in section A.2, our methods use a large VRAM for a long time in meta-training, a still large but relatively smaller VRAM for a very short time when feed forward through metanetwork during pruning, and a small VRAM for the standard finetuning.
+Specifically in this experiment, we use 10000 MiB for meta-training, 6000 MiB for feedforward through metanetwork, and 1000 MiB for finetuning.
+NVIDIA A100 with 80 GiB VRAM is enough for meta-training large models like ResNet50 and ViT-B-16.
+When we don’t have enough memory capacity, we can also change the batch size and use more time to make up for the lack of our memory capacity.
+We’ve also tried reducing the size of metanetwork and fit on NVIDIA A100 40G and the results seem not much influenced.
+All finetuning are standard finetuning and can be done on NVIDIA RTX 4090 with 24G VRAM.
+In all, during pruning only the feed forward through metanetwork requires large VRAM for a very short time and anything else is plain finetuning and requires small VRAM, during meta-training the VRAM is in a resonable range.
 Appendix B Framework implementation details
 B.1 Metanetwork(GNN) Architecture
-We build our metanetwork based on the message passing framework (Gilmer et al., 2017) and PNA (Corso et al., 2020) architecture.
+We build our metanetwork based on the message passing framework (MessagePassing) and PNA (PNA) architecture.
 Notation:
@@
 B.2.1 DepGraph and Torch-Pruning
-DepGraph (Fang et al., 2023) proposes a general sparsity regularization based structural pruning framework that can be applied to a wide range of neural network architectures, including CNNs, RNNs, GNNs, Transformers, etc.
+DepGraph (DepGraph) proposes a general sparsity regularization based structural pruning framework that can be applied to a wide range of neural network architectures, including CNNs, RNNs, GNNs, Transformers, etc.
 Alongside the paper, the authors released Torch-Pruning, a powerful Python library that enables efficient structural pruning for most modern architectures.
-In our work, the pruning criterion is designed based on the methodology of DepGraph (Fang et al., 2023), and all pruning operations are implemented using Torch-Pruning.
-Figure 7: A picture from (Fang et al., 2023). It shows how we group parameters together in structural pruning.
+In our work, the pruning criterion is designed based on the methodology of DepGraph (DepGraph), and all pruning operations are implemented using Torch-Pruning.
+Figure 9: A picture from (DepGraph). It shows how we group parameters together in structural pruning.
 B.2.2 Structural pruning
-In all our experiments, we only do structural pruning.
-Although we did not conduct experiments with unstructured pruning, our implementations are directly applicable to that setting as well.
+We perform most of our experiments by default using structural pruning, but also include a small number of experiments with unstructured and N:M sparsity pruning to demonstrate that our method can be applied to all kinds of pruning.
+Here we mainly introduce the structural pruning.
 Structural pruning involves removing parameters in predefined groups, so that the resulting pruned model can be used as a standalone network—without relying on the original model with masks, nor requiring specialized AI accelerators or software to realize reductions in memory footprint and computational cost.
@@
 We define the number of prunable dimensions of a group as the size of the input channel dimension of Conv $f_{4}$ (which is equivalent to the corresponding dimensions in BN $f_{2}$, the output channel of Conv $f_{1}$, etc.)
-For a more comprehensive theory to find all pruning groups in structural pruning, refer to the DepGraph (Fang et al., 2023) paper (section 3.1 & 3.2).
+For a more comprehensive theory to find all pruning groups in structural pruning, refer to the DepGraph (DepGraph) paper (section 3.1 & 3.2).
 B.2.3 A importance score function
@@
 B.2.6 Pay Attention
-Though we draw ideas from Fang et al.
-(2023) and use their Torch-Pruning libirary, we find their paper is inconsistent with their code.
+Though we draw ideas from DepGraph and use their Torch-Pruning libirary, we find their paper is slightly inconsistent with their code.
 We wrote appendix B.2.3 & B.2.4 strictly based on their code, which is the Torch-Pruning library.
-If you read the Fang et al.
-(2023) paper section 3.3, you may find our description a little different from theirs, because their paper is inconsistent with their code, and we are in line with their code.
+If you read the DepGraph paper section 3.3, you may find our description a little different from theirs, because their paper is slightly inconsistent with their code, and we are in line with their code.
+DepGraph is really a good paper.
+The small inconsistency may are made unintentionally by the author, not on purpose.
 Appendix C Experimental Preliminaries
@@
 We typically do not visualize all metanetworks but instead search for the best one using a binary search strategy.
+Relationship between "Acc VS.
+Speed Up Curve" and Pruning Performance: Empirically, pruning performance can be qualitatively predicted from the "Accuracy VS.
+Speed Up" curve.
+The accuracy in the flat region of the curve typically represents the maximum achievable accuracy; after pruning and finetuning, the final accuracy is usually the same or slightly lower, and sometimes only a little bit higher at most.
+The amount of speed up that preserves this top accuracy is generally larger than the speed up at the curve’s turning point, and these two values tend to correlate—the larger the turning point speed up, the larger the maintainable speed up.
+Therefore, if we aim for a pruned model with, for example, 93% accuracy, we should select a metanetwork whose flat-region accuracy is around or slightly above 93%, and whose flat region is as long as possible.
+We cannot give a teoretical guarantee between the "Acc VS.
+Speed Up" curve and the pruning performance.
+But this is not only our problem, because as far as we know, all other pruning methods can’t give theoretical gurarantee between their pruning criterion and final performance as well.
+The reasons can be the network itself is too complex and processes like finetuning changes everything in an unpredictable way.
 Big Batch Size and Small Batch Size: We use two different batch sizes in our experiments: a small batch size and a big batch size.
@@
 Method Base Pruned $\Delta$ Acc Pruned FLOPs Speed Up
-NISP (Yu et al., 2018) 43.2% 1.76
-Geometric (He et al., 2019b) 93.59 93.26 -0.33 41.2% 1.70
-Polar (Zhuang et al., 2020) 93.80 93.83 0.03 46.8% 1.88
-DCP-Adapt (Zhuang et al., 2018) 93.80 93.81 0.01 47.0% 1.89
-CP (Li et al., 2017a) 92.80 91.80 -1.00 50.0% 2.00
-AMC (He et al., 2018b) 92.80 91.90 -0.90 50.0% 2.00
-HRank (Lin et al., 2020a) 93.26 92.17 -1.09 50.0% 2.00
-SFP (He et al., 2018a) 93.59 93.36 -0.23 52.6% 2.11
-ResRep (Ding et al., 2021) 93.71 93.71 0.00 52.8% 2.12
-SCP (Kang & Han, 2020) 93.69 93.23 -0.46 51.5% 2.06
-FPGM (He et al., 2019a) 93.59 92.93 -0.66 52.6% 2.11
-FPC (He et al., 2020) 93.59 93.24 -0.35 52.9% 2.12
-DMC (Gao et al., 2020) 93.62 92.69 -0.93 50.0% 2.00
-GNN-RL (Yu et al., 2022) 93.49 93.59 0.10 54.0% 2.17
-DepGraph w/o SL (Fang et al., 2023) 93.53 93.46 -0.07 52.6% 2.11
-DepGraph with SL (Fang et al., 2023) 93.53 93.77 0.24 52.6% 2.11
-ATO (Wu et al., 2024) 93.50 93.74 0.24 55.0% 2.22
+NISP (NISP) 43.2% 1.76
+Geometric (Geometric) 93.59 93.26 -0.33 41.2% 1.70
+Polar (Polar) 93.80 93.83 0.03 46.8% 1.88
+DCP-Adapt (DCP-Adapt) 93.80 93.81 0.01 47.0% 1.89
+CP (CP) 92.80 91.80 -1.00 50.0% 2.00
+AMC (AMC) 92.80 91.90 -0.90 50.0% 2.00
+HRank (HRank) 93.26 92.17 -1.09 50.0% 2.00
+SFP (SFP) 93.59 93.36 -0.23 52.6% 2.11
+ResRep (ResRep) 93.71 93.71 0.00 52.8% 2.12
+SCP (SCP) 93.69 93.23 -0.46 51.5% 2.06
+FPGM (FPGM) 93.59 92.93 -0.66 52.6% 2.11
+FPC (FPC) 93.59 93.24 -0.35 52.9% 2.12
+DMC (DMC) 93.62 92.69 -0.93 50.0% 2.00
+GNN-RL (GNN-RL) 93.49 93.59 0.10 54.0% 2.17
+DepGraph w/o SL (DepGraph) 93.53 93.46 -0.07 52.6% 2.11
+DepGraph with SL (DepGraph) 93.53 93.77 0.24 52.6% 2.11
+ATO (ATO) 93.50 93.74 0.24 55.0% 2.22
 Meta-Pruning (ours) 93.51 93.64 0.13 56.5% 2.30
@@
 Meta-Pruning (ours) 93.51 93.78 0.27 56.5% 2.30
-GBN (You et al., 2019a) 93.10 92.77 -0.33 60.2% 2.51
-AFP (Ding et al., 2018) 93.93 92.94 -0.99 60.9% 2.56
-C-SGD (Ding et al., 2019) 93.39 93.44 0.05 60.8% 2.55
-Greg-1 (Wang et al., 2021) 93.36 93.18 -0.18 60.8% 2.55
-Greg-2 (Wang et al., 2021) 93.36 93.36 0.00 60.8% 2.55
-DepGraph w/o SL (Fang et al., 2023) 93.53 93.36 -0.17 60.2% 2.51
-DepGraph with SL (Fang et al., 2023) 93.53 93.64 0.11 61.1% 2.57
-ATO (Wu et al., 2024) 93.50 93.48 -0.02 65.3% 2.88
+GBN (GBN) 93.10 92.77 -0.33 60.2% 2.51
+AFP (AFP) 93.93 92.94 -0.99 60.9% 2.56
+C-SGD (C-SGD) 93.39 93.44 0.05 60.8% 2.55
+Greg-1 (Greg) 93.36 93.18 -0.18 60.8% 2.55
+Greg-2 (Greg) 93.36 93.36 0.00 60.8% 2.55
+DepGraph w/o SL (DepGraph) 93.53 93.36 -0.17 60.2% 2.51
+DepGraph with SL (DepGraph) 93.53 93.64 0.11 61.1% 2.57
+ATO (ATO) 93.50 93.48 -0.02 65.3% 2.88
 Meta-pruning (ours) 93.51 93.64 0.13 65.6% 2.91
@@
 Method Base Pruned $\Delta$ Acc Pruned FLOPs Speed Up
-OBD (Wang et al., 2019) 73.34 60.70 -12.64 82.55% 5.73
-OBD (Wang et al., 2019) 73.34 60.66 -12.68 83.58% 6.09
-EigenD (Wang et al., 2019) 73.34 65.18 -8.16 88.64% 8.80
-Greg-1 (Wang et al., 2021) 74.02 67.55 -6.67 88.69% 8.84
-Greg-2 (Wang et al., 2021) 74.02 67.75 -6.27 88.69% 8.84
-DepGraph w/o SL (Fang et al., 2023) 73.50 67.60 -5.44 88.73% 8.87
-DepGraph with SL (Fang et al., 2023) 73.50 70.39 -3.11 88.79% 8.92
+OBD (OBD) 73.34 60.70 -12.64 82.55% 5.73
+OBD (OBD) 73.34 60.66 -12.68 83.58% 6.09
+EigenD (OBD) 73.34 65.18 -8.16 88.64% 8.80
+Greg-1 (Greg) 74.02 67.55 -6.67 88.69% 8.84
+Greg-2 (Greg) 74.02 67.75 -6.27 88.69% 8.84
+DepGraph w/o SL (DepGraph) 73.50 67.60 -5.44 88.73% 8.87
+DepGraph with SL (DepGraph) 73.50 70.39 -3.11 88.79% 8.92
 Meta-Pruning (ours) 73.65 68.65 -5.00 88.81% 8.94
@@
 Method Base Top-1(Top-5) Pruned Top-1($\Delta$) Pruned Top-5($\Delta$) Pruned FLOPs
-DCP (Zhuang et al., 2018) 76.01%(92.93%) 74.95%(-1.06%) 92.32%(-0.61%) 55.6%
-CCP (Peng et al., 2019) 76.15%(92.87%) 75.21%(-0.94%) 92.42%(-0.45%) 54.1%
-FPGM (He et al., 2019a) 76.15%(92.87%) 74.83%(-1.32%) 92.32%(-0.55%) 53.5%
-ABCP (Lin et al., 2020b) 76.01%(92.96%) 73.86%(-2.15%) 91.69%(-1.27%) 54.3%
-DMC (Gao et al., 2020) 76.15%(92.87%) 75.35%(-0.80%) 92.49%(-0.38%) 55.0%
-Random (Li et al., 2022) 76.15%(92.87%) 75.13%(-1.02%) 92.52%(-0.35%) 51.0%
-DepGraph (Fang et al., 2023) 76.15%(-) 75.83%(-0.32%) - 51.7%
-ATO (Wu et al., 2024) 76.13%(92.86%) 76.59%(+0.46%) 93.24%(+0.38%) 55.2%
-DTP (Li et al., 2023) 76.13%(-) 75.55%(-0.58%) - 56.7%
+DCP (DCP-Adapt) 76.01%(92.93%) 74.95%(-1.06%) 92.32%(-0.61%) 55.6%
+CCP (CCP) 76.15%(92.87%) 75.21%(-0.94%) 92.42%(-0.45%) 54.1%
+FPGM (FPGM) 76.15%(92.87%) 74.83%(-1.32%) 92.32%(-0.55%) 53.5%
+ABCP (ABCP) 76.01%(92.96%) 73.86%(-2.15%) 91.69%(-1.27%) 54.3%
+DMC (DMC) 76.15%(92.87%) 75.35%(-0.80%) 92.49%(-0.38%) 55.0%
+Random (Random) 76.15%(92.87%) 75.13%(-1.02%) 92.52%(-0.35%) 51.0%
+DepGraph (DepGraph) 76.15%(-) 75.83%(-0.32%) - 51.7%
+ATO (ATO) 76.13%(92.86%) 76.59%(+0.46%) 93.24%(+0.38%) 55.2%
+DTP (DTP) 76.13%(-) 75.55%(-0.58%) - 56.7%
 ours 76.14%(93.11%) 76.13%(-0.01%) 92.78%(-0.33%) 57.2%
@@
 pruner reg 10
-Final Pruning metanetwork epoch 12
+Final Pruning metanetwork epoch 16
 speed up 2.3095
@@
 ResNet110 93.04 93.38 92.40
+F.3 Transfer From Small Dataset to Large Dataset
+When we are using large dataset like imagenet, is it possible that we use a much smaller dataset during meta-training but get the same results? Our answer is yes.
+When pruning Resnet50 on IMAGENET, we evenly choose 10% on each class in IMAGENET and form a new subset dataset.
+Meta-Training using this subset causes almost no drop in accuracy.
+See results in Table 19.
+This suggest that even 10% percent dataset is enough in meta-train to let metanetwork learn the pruning strategy.
+Table 19: During meta-train, use full ImageNet vs. only 10% ImageNet
+Method Base Top-1(Top-5) Pruned Top-1($\Delta$) Pruned Top-5($\Delta$) Pruned FLOPs
+ours 76.14%(93.11%) 76.13%(-0.01%) 92.78%(-0.33%) 57.2%
+ours(10%) 76.14%(93.11%) 76.24%(+0.10%) 92.65%(-0.46%) 57.0%
+Table 20: Transfer classification to detection
+Method Origin Prune w/o metanetwork Prune with metanetwork
+mAP 0.6061 0.4524 0.5173
 Appendix G Expriments on Transformers
@@
 All results are summarized in Table 21.
-Note that unlike CNNs, Vision Transformers naturally lacks inductive bias and need much more epochs of finetuning to reach higher accuracy.
-For other works like DepGraph (Fang et al., 2023), theirs method need to finetune 600 epochs in total, among them, 300 epochs are sparsity learning (slower than usual finetuning) and it takes more than a week to run a single round.
-We don’t have enough time to fully run the experiments with more finetuning for now, but we will continue tune our hyperparameters, add more epochs and hopefully give a better results in the next version of our paper.
 G.6 Hyperparameters
@@
 […]
-ViT-B/16 (Dosovitskiy et al., 2021) 81.07% - - 17.6
-CP-ViT (Song et al., 2022) 77.91% 77.36% -0.55% 11.7
-DepGraph (Fang et al., 2023) 81.07% 79.17% -1.90% 10.4
-MetaPruning(ours) 81.07% 77.59% -2.83% 11.0
+ViT-B/16 (ViT) 81.07% - - 17.6
+CP-ViT (Cp-vit) 77.91% 77.36% -0.55% 11.7
+DepGraph (DepGraph) 81.07% 79.17% -1.90% 10.4
+MetaPruning(ours) 81.07% 78.26% -2.81% 10.3
 Table 22: ViT-B/16 on ImageNet Hyperparameters
@@
 edge res ratio 0.1
-Meta Training lr 0.01
+Meta Training lr 0.001
 weight decay 0.0005
-milestone "3"
-pruner reg 1000
-Final Pruning metanetwork epoch 6
-speed up 1.6000
-Finetuning After Metanetwork epoch 100
-lr 0.0001
+milestone -
+pruner reg 10
+Final Pruning metanetwork epoch 22
+speed up 1.7
+Finetuning After Metanetwork epoch 300
+lr 0.01
 weight decay 0.01
 scheduler cosineannealinglr
+label smoothing 0.1
+mixup alpha 0.2
+cutmix alpha 0.1
 Finetuning After Pruning epoch 100
@@
 scheduler cosineannealinglr
-Appendix H The Use of Large Language Models
+label smoothing 0.1
+mixup alpha 0.2
+cutmix alpha 0.1
+Appendix H More Visualization of Statistics
+Figure 13: $l_{1}$ Norm
+Figure 14: Inter-Channel Correlation
+Figure 15: Effeicient Rank
+Appendix I The Use of Large Language Models
 We only use LLMs to check grammar errors and polish our writing. All other works are done by human writers.